### v2

# Plain Transformers Can be Powerful Graph Learners

Abstract
Transformers have attained outstanding performance across various modalities, owing to their simple but powerful scaled-dot-product (SDP) attention mechanisms. Researchers have attempted to migrate Transformers to graph learning, but most advanced Graph Transformers (GTs) have strayed far from plain Transformers, exhibiting major architectural differences either by integrating message-passing or incorporating sophisticated attention mechanisms. These divergences hinder the easy adoption of training advances for Transformers developed in other domains. Contrary to previous GTs, this work demonstrates that the plain Transformer architecture can be a powerful graph learner. To achieve this, we propose to incorporate three simple, minimal, and easy-to-implement modifications to the plain Transformer architecture to construct our Powerful Plain Graph Transformers (PPGT): (1) simplified $L_{2}$ attention for measuring the magnitude closeness among tokens; (2) adaptive root-mean-square normalization to preserve token magnitude information; and (3) a simple MLP-based stem for graph positional encoding. Consistent with its theoretical expressivity, PPGT demonstrates noteworthy realized expressivity on the empirical graph expressivity benchmark, comparing favorably to more complicated competitors such as subgraph GNNs and higher-order GNNs. Its outstanding empirical performance across various graph datasets also justifies the practical effectiveness of PPGT.
1 Introduction
Transformers have achieved excellent performance across various domains, from language Vaswani et al., (2017); Devlin et al., (2019); Brown et al., (2020) to vision Dosovitskiy et al., (2021); Touvron et al., 2021a ; Touvron et al., 2021b , and are well known for their reduced dependency on inductive bias as well as stronger flexibility and generalizability Dosovitskiy et al., (2021). The recent success of transformer-based, multi-modal, large language models (LLMs) OpenAI, (2024); Dubey et al., (2024) has also brought the once unattainable dream of building artificial general intelligence closer to reality. However, these modalities are primarily data on Euclidean spaces, while non-Euclidean domains are still under-explored. Non-Euclidean domains like graphs are important for describing physical systems with different symmetry properties, from protein structure prediction Jumper et al., (2021) to complex physics simulation Sanchez-Gonzalez et al., (2018). Therefore, many researchers have attempted to migrate Transformer architectures for graph learning for the future integration into multimodal foundation models.
Nevertheless, unlike other Euclidean domains, the characteristics of graphs make the naive migration challenging. Following the suboptimal performance of an early attempt Dwivedi and Bresson, (2021), researchers introduced significant architectural modifications to make Transformers perform well for graph learning. These include: implicit/explicit message-passing mechanisms Kreuzer et al., (2021); Rampášek et al., (2022), edge-updating Kim et al., (2022); Hussain et al., (2022); Ma et al., (2023), and sophisticated non-SDP (scaled-dot-product) attention designs Chen et al., (2022); Ma et al., (2023). These complexities hinder the integration of advances for Transformers arising from other domains and impede progress towards potential future unification of multi-modalities.
In this work, we step back and rethink the difficulties that a plain Transformer architecture faces when processing graph data. Instead of adding significant architectural modifications, we propose three minimalist but effective modifications to empower plain Transformers with advanced capabilities for learning on graphs (as shown in Fig. 1(a)). This results in our proposed Powerful Plain Graph Transformers (PPGTs), which have remarkable empirical and theoretical expressivity for distinguishing graph structures and superior empirical performance on real-world graph datasets. These modifications can be incorporated without significantly changing the architecture of the plain Transformer. Thus, our proposal both retains the simplicity and generality of the Transformer and offers the potential of facilitating cross-modality unification.
((a)) The Macro Archit.
((b)) GraphGPS Block
((c)) GRIT Block
((d)) PPGT Block
[…]
2 Preliminaries
2.1 Graph Learning
Learning of Graphs and Encoding of Multisets
[…]
Graph Transformers – Learning Graphs with Pseudo-coordinates
The philosophy of Graph Transformers (GTs) is to learn graph representations using positional encodings (PEs) rather than operating directly on the original input graphs. Zhang et al., 2023b provide a theoretical framework (GD-WL) that demonstrates that GTs’ stronger expressivity stems from finer-grained information beyond 1-WL encoded in graph PEs. Zhang et al., 2024b demonstrate that GD-WL not only applies to distance-like or relative graph PEs but also enables analysis of absolute PEs (e.g., Laplacian PE) by transforming them into their relative counterparts. Concurrently, Ma et al., (2024) interpret graph PE as pseudo-coordinates in graph spaces that mimic manifolds, demonstrating that it can extend beyond graph distance and is not restricted to the Transformer architecture (e.g., convolutions on pseudo-coordinates). We follow the same philosophy and show that with a simple stem on pseudo-coordinates, even plain Transformers can achieve outstanding expressive power in graph learning.
2.2 Limitations in Plain Transformer Architectures
The Loss of Magnitude Information in Token-wise Normalization Layer
LayerNorm (LN Ba et al., (2016)) and Root-Mean-Square-Norm (RMSN Zhang and Sennrich, (2019)) are two widely used token-wise normalization techniques in Transformer-based models that effectively control token magnitudes:
$\text{LN}(\mathbf{x})=\frac{\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}\|}\odot\boldsymbol{\gamma}+\boldsymbol{\beta}\,,\quad$ $\text{RMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\odot\boldsymbol{\gamma}\,.$ (1)
[…]
Pitfalls of Scaled Dot-product Attention
With a good balance of capacity and efficiency, Scaled dot-product (SDP) attention has become the most foundational attention mechanism in modern Transformers Vaswani et al., (2017). It has been widely explored in previous works and well optimized in deep learning libraries.
However, SDP attention is not perfect. For query and key tokens $\mathbf{q}_{i},\mathbf{k}_{j}\in\mathbb{R}^{D}$, SDP attention is:
$\alpha_{ij}:=\mathtt{Softmax}_{j}(\hat{\alpha}_{ij})=\frac{\exp(\hat{\alpha}_{ij})}{\sum_{j^{\prime}}\exp(\hat{\alpha}_{ij^{\prime}})}\,,\text{where }\hat{\alpha}_{ij}:=\frac{\mathbf{q}_{i}^{\intercal}\mathbf{k}_{j}}{\sqrt{D}}=\frac{\cos(\mathbf{q}_{i},\mathbf{k}_{j})\cdot\|\mathbf{q}_{i}\|\cdot\|\mathbf{k}_{j}\|}{\sqrt{D}}\,.$ (2)
Here $\cos(\mathbf{q}_{i},\mathbf{k}_{j})\in[-1,1]$ is the cosine similarity, measuring the angle between $\mathbf{q}_{i}$ and $\mathbf{k}_{j}$, independent of the vector magnitudes.
Figure 2: Illustration for comparing different attention mechanisms: (b) visualization of attention scores. SDP attention is biased towards larger-magnitude $k_{2}$. Cos attention disregards the magnitude information. $L_{2}$ attention strikes a balance among SDP and Cos attention to attend to $k_{1}$, which has the lowest $L_{2}$ distance to the query $q$.
Three drawbacks of SDP attention necessitate the additional control of token magnitudes:
  1. 1.
    Softmax saturation: large token magnitudes (i.e., $\|\mathbf{q}_{i}\|$, $\|\mathbf{k}_{j}\|$) lead to large pre-softmax-logit values $\hat{\alpha}_{ij}$ and consequently extremely small gradients in backpropagation Vaswani et al., (2017).
  2. 2.
    No closeness measurement on magnitude: for each query, $\|\mathbf{q}_{i}\|$ degenerates to a temperature factor $\tau$, and only controls the sharpness of attention scores, irrespective of the closeness between $\|\mathbf{q}_{i}\|$ and $\|\mathbf{k}_{j}\|$ (Fig.  2).
  3. 3.
    Biased to large magnitude keys: with $\|\mathbf{q}_{i}\|$ as the temperature $\tau$, compared to $\cos(\mathbf{q}_{i},\mathbf{k}_{j})\in[-1,1]$, large $\|\mathbf{k}_{j}\|\in[0,+\infty)$ will dominate attention scores.
To mitigate the first and third drawbacks, existing plain Transformers heavily rely on token-wise normalization to regulate the token magnitudes. Removing LN/RMSN and/or using BatchNorm(BN) Ioffe and Szegedy, (2015) instead can induce training instability and divergence Touvron et al., 2021b ; Yao et al., (2021).
3 Method
As established in Section 2, plain Transformers exhibit suboptimal performance in graph learning, due to the architectural limitations that hinder preserving and modeling token magnitude information — a critical aspect of capturing graph structures. Consequently, previous GTs attempt to address it by adopting BN and overly sophisticated attention designs (e.g., MPNNs and/or MLP-based attention) Kreuzer et al., (2021); Rampášek et al., (2022); Ma et al., (2023). They have thus strayed away from plain Transformers, impeding the transfer of previously explored training advances and the potential unification of other foundation models.
In this work, we propose two minimal and easy-to-adapt modifications to directly address the aforementioned limitations in plain Transformers. Furthermore, we introduce an extra enhancement in the PE stem to boost the information extraction of PE. These enhancements enable plain Transformers to achieve outstanding expressive power for graph learning.
3.1 AdaRMSN
As discussed above, SDP attention mechanisms necessitate additional control on token magnitudes. However, existing token-wise normalization layers lead to irreversible information loss on magnitudes. We desire a normalization layer that can not only control the token magnitudes, but is also capable of preserving the magnitude information when necessary.
Inspired by adaptive normalization layers Dumoulin et al., (2017); de Vries et al., (2017); Peebles and Xie, (2023), we propose adaptive RMSN (AdaRMSN)
$\text{AdaRMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\cdot\gamma^{\prime}(\mathbf{x})\,,\quad\text{where }\gamma^{\prime}(\mathbf{x}):=\frac{1}{\sqrt{D}}\|\boldsymbol{\alpha}\cdot\mathbf{x}+\boldsymbol{\beta}\|\,.$ (3)
The parameter $\boldsymbol{\beta}\in\mathbb{R}^{D}$ is initialized as $\mathbf{1}$ and $\boldsymbol{\alpha}\in\mathbb{R}^{D}$ as $\mathbf{0}$, leading to $\gamma^{\prime}(\mathbf{x})=1$. AdaRMSN behaves the same as regular RMSN at the initial stage of training, but is capable of recovering the identity transformation with $\boldsymbol{\beta}=\mathbf{0}$ and $\boldsymbol{\alpha}=\mathbf{1}$ when necessary.
3.2 Simplified $L_{2}$ Attention Mechanisms
From Dot-product to Euclidean Distance
To empower SDP attention to sense both angle- and magnitude-information among query and key tokens, we revisit $L_{2}$ attention Kim et al., (2021), which is based on Euclidean distance and is capable of measuring token closeness by balancing both angles and magnitudes, in contrast to SDP attention and cosine-similarity attention (as shown in Fig. 2).
To achieve better alignment with SDP attention, we further simplify it and reformulate it as SDP-attention with an additional bias term:
$\alpha_{ij}$ $:=\mathtt{Softmax}_{j}(-\frac{1}{\sqrt{D}}\cdot\frac{1}{2}\|\mathbf{q}_{i}-\mathbf{k}_{j}\|_{2}^{2})=\mathtt{Softmax}_{j}(\frac{1}{\sqrt{D}}(\mathbf{q}_{i}^{\intercal}\mathbf{k}_{j}-\frac{1}{2}\mathbf{q}_{i}^{\intercal}\mathbf{q}_{i}-\frac{1}{2}\mathbf{k}_{j}^{\intercal}\mathbf{k}_{j}))$
$=\mathtt{Softmax}_{j}(\frac{1}{\sqrt{D}}\mathbf{q}_{i}^{\intercal}\mathbf{k}_{j}-\frac{1}{2\sqrt{D}}\mathbf{k}_{j}^{\intercal}\mathbf{k}_{j})\,.$ (4)
This can be easily supported by existing SDP attention implementations in most deep learning libraries, e.g., PyTorch Paszke et al., (2019) 111e.g., the bias term can be directly formulated as a float attention mask in SDP-attention in PyTorch..
[…]
Let $\mathbf{p}_{ij}$ denote the relative positional embeddings for node-pair $(i,j)$ shared by all attention blocks, which is potentially processed by the stems (preprocessing modules). For each head, the attention scores $\alpha_{ij}\in\mathbb{R}$ for query/key tokens $\mathbf{q}_{i},\mathbf{k}_{j}\in\mathbb{R}^{D}$ in the proposed attention are computed as
$\alpha_{ij}$ $:=\phi(\mathbf{p}_{ij})\cdot\mathtt{Softmax}_{j}\big{(}\frac{\mathbf{q}_{i}^{\intercal}\mathbf{k}_{j}}{\sqrt{D}}-\frac{\mathbf{k}_{j}^{\intercal}\mathbf{k}_{j}}{2\sqrt{D}}+\theta(\mathbf{p}_{ij})\big{)}\,,$ (5)
where $\phi:\mathbb{R}^{D}\to\mathbb{R}$ and $\theta:\mathbb{R}^{D}\to\mathbb{R}$ are linear transforms. The $\phi(\mathbf{\mathbf{p}_{ij}})$ is an optional term purely based on the relative position to guarantee the universality of attention with relative PE (URPE) Luo et al., (2022). This term demands slight customization of the existing attention implementation, but we retain it since it is beneficial for learning objects with complicated structures Luo et al., (2022); Zhang et al., 2023b .
Note that, the attention with PE in relative-form is general, since it is widely employed in many existing Transformers, from language Shaw et al., (2018); Raffel et al., (2020); Press et al., (2022) to vision Dosovitskiy et al., (2021); Liu et al., (2021, 2022). Many absolute PEs are de facto explictly/implictly transformed into the relative-form in attention mechanisms Su et al., (2024); Huang et al., (2024); Zhang et al., 2024b .
3.3 Powerful Plain Graph Transformers
In this section, we provide an example of constructing plain Transformers for learning graphs with our proposed techniques, termed Powerful Plain Graph Transformers (PPGT).
We follow the plain Vision Transformers architecture Dosovitskiy et al., (2021); Touvron et al., 2021a , — stems, backbone and prediction head (as shown in Fig. 1(a)(a) and Fig. 1(a)(d)).
Graph Positional Encoding
In this work, we utilize relative random walk probabilities (RRWP) Ma et al., (2023) as our demonstrating example of graph PE, considering its simplicity and effectiveness. RRWP is defined as
$\mathbf{p}^{\prime}_{ij}=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\dots,\mathbf{M}^{K-1}]_{[i,j]}\in\mathbb{R}^{K}\,,$ (6)
where $\mathbf{X}_{[i,j]}$ stands for the $i,j$th element/slice of a tensor $\mathbf{X}$; $\mathbf{M}:=\mathbf{D}^{-1}\mathbf{A}$ is the random walk matrix given the adjacency matrix $\mathbf{A}$ of the graph; and $\mathbf{I}\in\mathbb{R}^{N\times N}$ denotes the identity matrix. Most other graph PEs are applicable in our framework with minor modifications to the PE stem.
Transformer Backbone
Following the latest plain Transformers, we utilize the pre-norm Xiong et al., (2020) architecture, for $l=1,\cdots,L$,
$\hat{\mathbf{X}}^{l}=\mathbf{X}^{l-1}+\mathtt{MSA}(\mathtt{Norm}(\mathbf{X}^{l-1}),\mathbf{P}),\quad\mathbf{X}^{l}=\mathtt{FFN}(\mathbf{X}^{l}):=\hat{\mathbf{X}}^{l}+\mathtt{MLP}(\mathtt{Norm}(\hat{\mathbf{X}}^{l})),\quad\mathbf{Y}=\mathtt{Norm}(\mathbf{X}^{L})$ (7)
where $\mathbf{X}^{l}=[\mathbf{x}_{i}^{l}]_{i=1}^{N}\in\mathbb{R}^{N\times D}$ contains the node representations at layer $l$; $\mathbf{P}=[[\mathbf{p}_{ij}]_{i=1}^{N}]_{j=1}^{N}\in\mathbb{R}^{N\times N\times D}$ contains the relative positional embeddings; $\mathtt{MSA}$ denotes multihead self-attention; $\mathtt{Norm}$ indicates the normalization layer; $\mathtt{MLP}$ is a 2-layer multilayer perception and $\mathtt{FFN}$ is a Feedfoward networks with pre-norm.
Stem
Before the Transformer backbone, we use small networks, usually referred to as stems, to process positional encoding $\mathbf{p}^{\prime}_{ij}$ and merge node/edge attributes $\mathbf{x}^{\prime}_{i}$/$\mathbf{e}^{\prime}_{ij}$.
We consider a simple stem design for node and PE, respectively:
$\mathbf{x}^{0}_{i}=\mathtt{FC}(\mathbf{x}^{\prime}_{i})+\mathtt{FC}(\mathbf{p}^{\prime}_{ii}),\quad\mathbf{p}^{0}_{ij}=\mathtt{Norm}\circ\mathtt{FFN}\circ\cdots\circ\mathtt{FFN}(\mathtt{FC}(\mathbf{e}^{\prime}_{ij})+\mathtt{FC}(\mathbf{p}^{\prime}_{ij})),$ (8)
where $\mathtt{FC}$ stands for the fully-connected layer (e.g., linear projection); $\circ$ stands for function composition; $\mathbf{x}^{\prime}_{i}$ is treated as zero if there are no node attributes; $\mathbf{e}^{\prime}_{ij}$ is set to zero if there are no edge attributes or if $(i,j)$ is not an observed edge. Driven by the analysis of RRWP in Ma et al., (2023), we introduce additional FFNs and a final normalization layer in the PE stem to better extract the structural information, mimicking the pre-norm architecture of the Transformer backbone.
Prediction Head
We use a task-specific MLP prediction head, following GRIT Ma et al., (2023) and GraphGPS Rampášek et al., (2022): sum/mean-pooling followed by a MLP for graph level tasks; and a MLP shared by all nodes for node level tasks.
3.4 Sinusoidal PE Enhancement (SPE)
Figure 3: Illustration of two node-pairs (a) $(i,j)$ and $(i,k)$ of a graph, and (b) absolute difference of RRWPs and sinusoidally-encoded RRWPs for those two node-pairs.
As discussed in Zhang et al., 2023b ; Zhang et al., 2024b , the finer-grained the information provided through graph positional encodings (PEs), the stronger distinguishability of graph structure GD-GNNs can achieve. However, due to the spectral bias of neural networks Rahaman et al., (2019), MLPs prioritize learning the low-frequency modes and thus lose detailed information stored in PE.
To mitigate this issue, motivated by NeRF Mildenhall et al., (2020), we add an extra sinusoidal encoding on top of the RRWP and process it with a simple MLP. The sinusoidal encoding is applied to each channel of $\mathbf{p}^{\prime}_{ij}$ in an elementwise fashion:
$\mathtt{SPE}(p^{\prime}_{ijk})$ $=\big{[}p^{\prime}_{ijk},\sin(2^{0}\pi p^{\prime}_{ijk}),\ \cos(2^{0}\pi p^{\prime}_{ijk}),\dots,\quad\sin\big{(}2^{S-1}\pi p^{\prime}_{ijk}\big{)},\ \cos\big{(}2^{S-1}\pi p^{\prime}_{ijk}\big{)}\big{]}\,,$ (9)
where $\mathtt{SPE}:\mathbb{R}\to\mathbb{R}^{1+2S}$ with $S\in\mathbb{Z}^{+}$ different bases. For notational conciseness, we use $\mathtt{SPE}(\mathbf{p}^{\prime}_{ij})\in\mathbb{R}^{K+2SK}$ to denote the application of $\mathtt{SPE}$ to all channels of $\mathbf{p}^{\prime}_{ij}$ and concatenate the outputs. The PE stem (Eq. 8) becomes
$\mathbf{p}^{0}_{ij}=\mathtt{Norm}\circ\mathtt{FFN}\circ\cdots\circ\mathtt{FFN}(\mathtt{FC}(\mathbf{e}^{\prime}_{ij})+\mathtt{MLP}(\mathtt{SPE}(\mathbf{p}^{\prime}_{ij})))$ (10)
As shown in Fig. 3, after sinusoidal encoding, the signal differences between $\mathbf{p}^{\prime}_{ij}$ and $\mathbf{p}^{\prime}_{ik}$ are amplified.
3.5 Theoretical Expressivity of PPGT
The theoretical expressivity of PPGT can be analyzed with the GD-WL framework Zhang et al., 2023b . Since it is not proposed in our work, we provide further discussion of it in Appx. E.1.
4 Relationship with Previous Work
Graph Transformers Motivated by the success of Transformers in other domains, researchers have strived to migrate Transformers to graph learning to address the limitations of MPNNs. Although the naive migration of Transformer to graph learning did not work well Dwivedi and Bresson, (2021), several recent works have achieved considerable success when constructing graph Transformers, from theoretical expressivity analysis Zhang et al., 2023b ; Zhang et al., 2024b to outstanding empirical performance Ying et al., (2021); Rampášek et al., (2022); Ma et al., (2023). However, during the development of these graph Transformers, there has been a gradual but substantial deviation from the plain Transformers widely used in other domains. For example, SAN Kreuzer et al., (2021) introduces dual-attention mechanisms with local and global aggregations; K-Subgraph SAT Chen et al., (2022) introduce MPNNs into attention mechanisms to compute attention scores; GraphGPS Rampášek et al., (2022) heavily relies on the MPNNs within its hybrid Transformer architecture; EGT Hussain et al., (2022) introduces gated mechanisms and edge-updates inside the attention; GRIT Ma et al., (2023) incorporates a complicated conditional MLP-based attention mechanism (shown in Appendix. D.1). These deviations prevent the easy adoption of Transformer training advances and obscure the potential unification of cross-modality foundation models. Among these graph Transformers, the graphormer-series Ying et al., (2021); Luo et al., (2022); Zhang et al., 2023b retain architectures that are closest to plain Transformers; unfortunately, the result is a substantial gap in empirical performance and empirical expressivity to the best-performing graph Transformers Rampášek et al., (2022); Ma et al., (2023). In this work, we analyze the limitations of plain Transformers for learning on graphs and propose simple but effective designs to allow plain transformers to reach outstanding empirical performance as well as remarkable expressivity.
Other Attention Mechanisms Besides SDP attention, there are other attention variants in use. Bahdanau et al., (2015) introduced the earliest content-based attention mechanism for recurrent neural networks (RNNs) based on an MLP, which is more computational and memory costly. Swin-Transformer-V2 Liu et al., (2022) uses cosine-similarity to compute attention scores for better stability, but neglects the magnitude information. Kim et al., (2021) propose the use of the negative of the square $L_{2}$-distance, with tied query-key projection weights, for maintaining Lipchitz continuity of Transformers. Our attention mechanism, although based on $L_{2}$ attention, is further simplified and adjusted in order to maintain alignment with SDP attention.
Additional related work We discuss additional related work, including MPNNs, graph positional/structural encoding, higher-order GNNs, and subgraph GNNs in Appx. D.
5 Experimental Results
5.1 Empirical Expressivity on Graph Isomorphism
To better understand the expressivity of PPGT, we evaluate our model on the BREC benchmark Wang and Zhang, (2024), a comprehensive dataset for measuring the empirical expressive power of GNNs w.r.t. graph isomorphism, with graph-pairs from 1-WL to 4-WL-indistinguishable.
From the results (as shown in Tab. 1), we can uncover several conclusions and insights:
[1]. GTs can reach empirical expressivity approaching the theoretical expressivity: With proper architectural designs and graph PE, most GTs achieve decent expressivity bounded by 3-WL, matching 3-WL equivalence on the Basic, Regular, and Extended categories of graph pairs. Graphormer, as an example of earlier plain Transformers, demonstrates inferior expressivity, highlighting the necessity of improving plain Transformers for graph structure learning. PPGT, while maintaining a plain Transformer architecture, achieves powerful empirical expressivity through our proposed modifications, surpassing other GTs with more sophisticated architectures.
[2]. Mismatch between theoretical and empirical expressivity: The theoretical expressivity is not completely reflected in the empirical expressivity. For example, despite the same theoretical expressivity, adding SPE—which enhances the information extraction from PE—to PPGT can significantly boost the empirical expressivity, distinguishing 24 pairs of graphs in CFI (improved from 8 pairs). On the other hand, EPNN and PPGN, despite having stronger theoretical expressivity, achieve worse empirical expressivity compared to PPGT. This indicates that besides theoretical expressivity, whether GNNs can effectively learn to fulfill their theoretical expressivity also matters.
[3]. Mismatch between expressivity and real-world benchmark performance: The stronger theoretical/empirical expressivity is not completely reflected in real-world benchmark performance. For example, subgraph GNNs and/or K-WL GNNs with stronger expressivity (e.g., SSWL+, I2GNN, N2GNN) reach inferior performance compared to GRIT and PPGT on the ZINC benchmark.
[4]. Going beyond GD-WL?: As previously discussed, the design of graph PE, a.k.a., pseudo-coordiantes, can be extended beyond the distance/affinity of graphs. With this in mind, we conduct a exploratory demo called I2GNN+PPGT, which uses I2GNN to generate additional positional encodings for PPGT. The empirical expressivity is further improved to 76%, outperforming the standalone I2GNN and PPGT, surpassing 3-WL and reaching the top performance among the methods compared. This demonstration hints that plain Graph Transformers can potentially surpass GD-WL and achieve greater expressive power purely through enhanced positional encoding designs.
[…]
Basic (60) Reg.(140) Ext. (100) CFI (100) Total (400) ZINC
Type Model WL-Class Num.($\uparrow$) Num.($\uparrow$) Num.($\uparrow$) Num.($\uparrow$) Num.($\uparrow$) Acc. ($\uparrow$) MAE ($\downarrow$)
Heuristic Algorithm 1-WL 1-WL 0 0 0 0 0 0% -
3-WL 3-WL 60 50 100 60 270 67.5% -
[…]
K-WL GNNs PPGN 3-WL 60 50 100 23 233 58.2% -
[…]
I2GNN+PPGT GD+WL? 60 120 100 24 304 76% -
5.2 Benchmarking PPGT on Real-world Benchmarks
The specifications, details, and references for the baseline methods are provided in Appx. A.5.
Benchmarking GNNs
We conduct a general evaluation of our proposed PPGT on five datasets from Benchmarking GNNs (Dwivedi et al., 2022a, ): ZINC, MNIST, CIFAR10, PATTERN, and CLUSTER, and summarize the results in Table 2. We observe that our model obtains the best mean performance for all five datasets, outperforming various MPNNs, non-MPNN GNNs, and existing graph Transformers. These results showcase the effectiveness of PPGT for general graph learning with a plain Transformer architecture.
Table 2: Test performance in five benchmarks from Benchmarking GNNs (baselines please see Appx. A.5 for details and references). Shown is the mean $\pm$ s.d. of 4 runs with different random seeds. Highlighted are the top first, second, and third results. # Param under $500K$ for ZINC, PATTERN, CLUSTER and $\sim 100K$ for MNIST and CIFAR10.
Model ZINC MNIST CIFAR10 PATTERN CLUSTER
[…]
Message Passing Networks
GCN $0.367\pm 0.011$ $90.705\pm 0.218$ $55.710\pm 0.381$ $71.892\pm 0.334$ $68.498\pm 0.976$
GIN $0.526\pm 0.051$ $96.485\pm 0.252$ $55.255\pm 1.527$ $85.387\pm 0.136$ $64.716\pm 1.553$
GAT $0.384\pm 0.007$ $95.535\pm 0.205$ $64.223\pm 0.455$ $78.271\pm 0.186$ $70.587\pm 0.447$
GatedGCN $0.282\pm 0.015$ $97.340\pm 0.143$ $67.312\pm 0.311$ $85.568\pm 0.088$ $73.840\pm 0.326$
PNA $0.188\pm 0.004$ $97.94\pm 0.12$ $70.35\pm 0.63$- --
Non-MPNN Graph Neural Networks
CRaW1 $0.085\pm 0.004$ ${97.944}\pm{0.050}$ $69.013\pm 0.259$- --
GIN-AK+ ${0.080}\pm{0.001}$- $72.19\pm 0.13$ $86.850\pm 0.057$-
DGN $0.168\pm 0.003$- $\mathbf{72.838\pm 0.417}$ $86.680\pm 0.034$-
CKGCN $\mathbf{0.059\pm 0.003}$ $\mathbf{98.423\pm 0.155}$ $72.785\pm 0.436$ $\mathbf{88.661\pm 0.143}$ $79.003\pm 0.140$
Graph Transformers
SAN $0.139\pm 0.006$- -- $86.581\pm 0.037$ $76.691\pm 0.65$
K-Subgraph SAT $0.094\pm 0.008$- -- ${86.848\pm 0.037}$ $77.856\pm 0.104$
EGT $0.108\pm 0.009$ $\mathbf{98.173\pm 0.087}$ $68.702\pm 0.409$ $86.821\pm 0.020$ $\mathbf{79.232\pm 0.348}$
Graphormer-GD $0.081\pm 0.009$- -- -- --
GPS $0.070\pm 0.004$ ${98.051\pm 0.126}$ ${72.298\pm 0.356}$ $86.685\pm 0.059$ ${78.016\pm 0.180}$
GMLP-Mixer $0.077\pm 0.003$- -- -- --
GRIT $\mathbf{0.059\pm 0.002}$ $98.108\pm 0.111$ $\mathbf{76.468\pm 0.881}$ $\mathbf{87.196\pm 0.076}$ $\mathbf{80.026\pm 0.277}$
PGT (Ours) $\mathbf{0.0566\pm 0.002}$ $\mathbf{98.614\pm 0.096}$ $\mathbf{78.560\pm 0.700}$ $\mathbf{89.752\pm 0.030}$ $\mathbf{80.027\pm 0.114}$
Long Range Graph Benchmarks
We present experimental results on two Long-range Graph Benchmark (LRGB) Dwivedi et al., 2022c datasets, Peptides-Function and Peptides-Structure in Table 5.2. PPGT achieves the lowest MAE on Peptides-Structure and remains in the top three models on Peptides-Function. Moreover, the difference in AP for Peptides-Function dataset between the best performing GRIT and our PPGT is not statistically significant at the 5% level for a one-tailed t-test. These results demonstrate PPGT’s outstanding capability of learning long-range dependency structures.
PCQM4Mv2 Large-scale Graph Regression Benchmark
Further, we conduct an experiment on the PCQM4Mv2 large-scale graph regression benchmark [^2] Clarification on large-scale graph-datasets and large-scale-graph datasets can be found in Appx. F.1. of 3.7M graphs Hu et al., (2021), which is currently one of the largest molecular datasets (Table 5.2). We follow the experimental setting of Rampášek et al., (2022) — excluding 3D information as attributes, and treating the validation set of PCQM4Mv2 as the test set, instead of the original private Test-dev set. The result of a single random seed run is reported due to the size of the dataset, following previous works. We followed the hyperparameter configuration of GRIT Ma et al., (2023) and did not conduct hyperparameter search due to computational limits. However, our model can still reach an outstanding performance, comparable to GRIT and GraphGPS, with similar parameter budget.
Table 3: Test performance on two benchmarks from Long-range Graph Benchmarks (LRGB). (mean $\pm$ s.d. of 4 runs). Highlighted are the top first, second, and third results.
Model Peptides-func Peptides-struct
[…]
GCN $0.5930\pm 0.0023$ $0.3496\pm 0.0013$
GINE $0.5498\pm 0.0079$ $0.3547\pm 0.0045$
GatedGCN+RWSE $0.6069\pm 0.0035$ $0.3357\pm 0.0006$
CKGCN ${0.6952\pm 0.0068}$ ${0.2477\pm 0.0018}$
Transformer+LapPE $0.6326\pm 0.0126$ $0.2529\pm 0.0016$
SAN+LapPE $0.6384\pm 0.0121$ $0.2683\pm 0.0043$
SAN+RWSE ${0.6439\pm 0.0075}$ $0.2545\pm 0.0012$
GPS $0.6535\pm 0.0041$ $0.2500\pm 0.0012$
GMLP-Mixer $\mathbf{0.6970\pm 0.0080}$ $\mathbf{0.2475\pm 0.0015}$
GRIT $\mathbf{0.6988\pm 0.0082}$ $\mathbf{0.2460\pm 0.0012}$
PPGT (Ours) $\mathbf{0.6961\pm 0.0062}$ $\mathbf{0.2450\pm 0.0017}$
Table 4: Test performance on PCQM4Mv2 dataset. Shown is the result of a single run due to the computation constraint. Highlighted are the top first, second, and third results.
Method Model Valid. (MAE $\downarrow$) # Param
MPNNs GCN $0.1379$ 2.0M
GCN-virtual $0.1153$ 4.9M
GIN $0.1195$ 3.8M
GIN-virtual $0.1083$ 6.7M
GRPE $0.0890$ 46.2M
Graphormer $0.0864$ 48.3M
TokenGT (ORF) $0.0962$ 48.6M
Graph TokenGT (Lap) $0.0910$ 48.5M
Transformers GPS-small $0.0938$ 6.2M
GPS-medium $\mathbf{0.0858}$ 19.4M
GRIT $\mathbf{0.0859}$ 16.6M
PPGT (Ours) $\mathbf{0.0856}$ 17.6M
5.3 Ablation Study on Proposed Designs
Figure 4: Ablation Study on ZINC. MLPA: Conditional MLP Attention; DegS: degree scaler; ARMSN: AdaRMSNorm; URP: Universal RPE; SPE: Sinusoidal PE enhancement.
[…]
Further Study
We present further experimental studies and results in Appx. B.
6 Conclusion
Plain transformers are ill-suited for handling the unique challenges posed by graphs, such as the lack of canonical coordinates and permutation invariance. To obtain superior capacity and empirical performance, previous graph transformers (GTs) have introduced non-standard, sophisticated, and domain-specific architectural modifications. In this work, we demonstrate that plain transformers can be powerful graph learners via the proposed minimal, easy-to-adapt modifications. Our Powerful Plain Graph Transformers (PPGT) not only achieve competitive expressivity, but also demonstrate outstanding empirical performance on real-world graph benchmarks while maintaining the simplicity of plain Transformer architectures. Further exploration of empirical expressivity also unveils a potential direction for improving plain GTs. We consider this work an important first step toward reducing dissimilarities between GTs and transformers in other domains. This direction of investigation is crucial for building general multi-modal foundation models that integrate learning capabilities on graphs and potentially other irregular non-Euclidean geometric spaces.
Limitations:
This work demonstrates that plain Transformers can be powerful graph learners. However, like most plain Transformers, PPGTs still require the $O(N^{2})$ computational complexity and might encounter issues on scaling to large-size inputs. A more detailed discussion of limitations and possible avenues to address them is provided in Appx. F.
Appendix A Experimental Details
A.1 Bechmarking GNNs and Long Range Graph Benchmarks
Description of Datasets
Table 5 provides a summary of the statistics and characteristics of datasets used in this paper. The first five datasets are from Dwivedi et al., 2022a , and the last two are from Dwivedi et al., 2022c . Readers are referred to Rampášek et al., (2022) for more details of the datasets.
Table 5: Overview of the graph learning datasets involved in this work Dwivedi et al., 2022a ; Dwivedi et al., 2022c ; Irwin et al., (2012); Hu et al., (2021).
Dataset # Graphs Avg. # nodes Avg. # edges Directed Prediction level Prediction task Metric
ZINC 12,000 23.2 24.9 No graph regression Mean Abs. Error
SP-MNIST 70,000 70.6 564.5 Yes graph 10-class classif. Accuracy
SP-CIFAR10 60,000 117.6 941.1 Yes graph 10-class classif. Accuracy
PATTERN 14,000 118.9 3,039.3 No inductive node binary classif. Weighted Accuracy
CLUSTER 12,000 117.2 2,150.9 No inductive node 6-class classif. Weighted Accuracy
Peptides-func 15,535 150.9 307.3 No graph 10-task classif. Avg. Precision
Peptides-struct 15,535 150.9 307.3 No graph 11-task regression Mean Abs. Error
PCQM4Mv2 3,746,620 14.1 14.6 No graph regression Mean Abs. Error
Dataset splits, random seed and parameter budgets
We conduct the experiments on the standard train/validation/test splits of the evaluated benchmarks, following previous works Rampášek et al., (2022); Ma et al., (2023). For each dataset, we execute 4 trials with different random seeds (0, 1, 2, 3) and report the mean performance and standard deviation. We follow the most commonly used parameter budgets: around 500k parameters for ZINC, PATTERN, CLUSTER, Peptides-func, and Peptides-struct; and around 100k parameters for SP-MNIST and SP-CIFAR10.
Hyperparameters
Due to the limited time and computational resources, we did not perform an exhaustive search on the hyperparameters. We start with the hyperparameter setting of GRIT Ma et al., (2023) and perform minimal search to satisfy the commonly used parameter budgets.
The hyperparameters are presented in Table 6 and Table 7. In the tables, S.D. stands for stochastic depth Huang et al., (2016), a.k.a., drop-path, in which we treat a graph as one example in stochastic depth.
For Peptides-func, we find that there exists a mismatch between the cross-entropy loss and the metric, average precision, due to the highly imbalanced label distribution. Empirically, we observe that adding a BN in the prediction head post-graph-pooling effectively mitigates the problem.
Table 6: Hyperparameters for five datasets from Benchmarking GNNs Dwivedi et al., 2022a
Hyperparameter ZINC MNIST CIFAR10 PATTERN CLUSTER
PE Stem
MLP-dim 128 64 64 128 128
# FFN 2 1 2 2 2
FFN expansion 2 2 2 2 2
# bases 3 3 3 3 3
Backbone
# blocks 12 4 10 12 16
Dim 64 48 32 64 56
FFN expansion 2 2 2 2 2
# attn. heads 8 6 4 8 7
S.D. 0.1 0.2 0.2 0.1 0.3
Attn. dropout 0.2 0.5 0.2 0.1 0.3
Pred. Head
Graph Pooling sum mean mean - -
Norm - - - - -
# layers 3 2 2 2 2
PE-dim 24 10 10 32 32
Training
Batch size 32 32 16 32 16
Learn. rate 2e-3 1e-3 1e-3 1e-3 1e-3
# epochs 2500 400 200 400 150
# warmup 50 10 10 10 10
Weight decay 1e-5 1e-5 1e-5 1e-5 1e-5
# parameters 487K 102K 108K 497K 496K
Table 7: Hyperparameters for two datasets from the Long-range Graph Benchmark Dwivedi et al., 2022c
Hyperparameter Peptides-func Peptides-struct
PE Stem
MLP-dim 128 128
# FFN 1 1
FFN expansion 2 2
# bases 3 3
Backbone
# blocks 5 5
Dim 96 96
FFN expansion 2 2
# attn. heads 16 8
S.D. 0.1 0.1
Attn. dropout 0.2 0.1
Pred. Head
Graph Pooling sum mean
Norm BN -
# layers 3 2
PE-dim 32 24
Training
Batch size 32 32
Learn. rate 7e-4 7e-4
# epochs 400 250
# warmup 10 10
Weight decay 1e-5 1e-5
# Parameters 509K 488K
A.2 BREC: Empirical Expressivity Benchmark
BREC Wang and Zhang, (2024) is an empirical expressivity datasets, consisting of four major categories of graphs: Basic, Regular, Extension, and CFI. Basic graphs include $60$ pairs of simple 1-WL-indistinguishable graphs. Regular graphs include $140$ pairs of regular graphs from four types of subcategories: $50$ pairs of simple regular graphs, $50$ pairs of strongly regular graphs, $20$ pairs of 4-vertex condition graphs and $20$ pairs of distance regular graphs. Extension graphs include $100$ pairs of special graphs that arise when comparing four kinds of GNN extensions Papp and Wattenhofer, (2022). CFI graphs include $100$ pairs of graphs generated by CFI methods Cai et al., (1992). All pairs of graphs are 1-WL-indistinguishable.
We follow the standard training pipelines from BREC: pairwise contrastive training and evaluate process. No specific parameter budget required on BREC dataset.
Hyperparameter
The hyperparameter setting for PPGT on BREC can be found in Table 8.
For the experimental setup of I2-GNN + PPGT, we construct a subgraph for each edge within the 4-hop neighborhood and use RRWP as the node-attributes. We employ a 6-layer GIN with BN on each subgraph independently to get its representation. The subgraph representations are fed to PPGT as the additional edge-attributes. The I2-GNN and PPGT are trained end-to-end together.
Table 8: Hyperparameters for BREC Wang and Zhang, (2024)
Hyperparameter BREC
PE Stem
MLP-dim 192
# FFN 4
FFN expansion 2
# bases 15
Backbone
# blocks 6
Dim 96
FFN expansion 2
# attn. heads 16
S.D. 0.
Attn. dropout 0.
Pred. Head
Graph Pooling sum
Norm BN -
# layers 3
PE-dim 32
Training
Batch size 32
Learn. rate 1e-3
# epochs 200
# warmup 10
Weight decay 1e-5
# parameters 874K
A.3 PCQM4Mv2 from OGB Large-Scale Challenge
PCQM4Mv2 is a large-scale quantum chemistry graph dataset benchmark, containing over 3.7M graphs, proposed from OGB Large-Scale Challenge (OGB-LSC) Hu et al., (2021). The statistics of the dataset can be found in Table 5.
A.4 Optimizer and Learning Rate Scheduler
Following most plain Transformers in other domains, we use AdamW Loshchilov and Hutter, (2019) as the optimizer and the Cosine Annealing Learning Rate scheduler Loshchilov and Hutter, (2017) with linear warm up.
A.5 Baselines Information
For benchmarks from BenchmarkingGNN Dwivedi et al., 2022a .
  • Tyical Message-passing Networks(MPNNs): GCN Kipf and Welling, (2017), GIN Xu et al., (2019), GAT Veličković et al., (2018), GatedGCN Bresson and Laurent, (2018), PNA Corso et al., (2020);
  • GNNs going beyond MPNNs: CRaW1 Tönshoff et al., (2023), GIN-AK+ Zhao et al., (2022), DGN Beani et al., (2021), CKGCN Ma et al., (2024),
  • Graph Transformers SAN Kreuzer et al., (2021), K-Subgraph SAT Chen et al., (2022), EGT Hussain et al., (2022), Graphormer-GD Zhang et al., 2023b , GPS Rampášek et al., (2022), GMLP-Mixer He et al., (2023), GRIT Ma et al., (2023).
For BREC Wang and Zhang, (2024):
  • Subgraph GNNs: SUN Frasca et al., (2022), SSWL+ Zhang et al., 2023a , I2-GNN Huang et al., (2023)
  • K-WL/K-FWL GNNs: PPGN Maron et al., (2019), 2-DRFWL(2) Zhou et al., (2023), 3-DRFWL(2) Zhou et al., (2023)
  • K-FWL+SubgraphGNNs: N2GNN Feng et al., (2023)
  • GD-WL GNNs Graphormer Ying et al., (2021), EPNN Zhang et al., 2024b , CKGConv Ma et al., (2024), GRIT Ma et al., (2023).
For OGB-LSC Hu et al., (2021):
MPNNs (GCN Kipf and Welling, (2017), GIN Xu et al., (2019) with/without virtual nodes) as well as several Graph Transformers (GRPE Park et al., (2022), Graphormer Ying et al., (2021), TokenGT Kim et al., (2022) and GraphGPS Rampášek et al., (2022)).
Appendix B Additional Study
B.1 Sensitivity Study: Impact of the Number of Bases $S$ in SPE on Empirical Expressivity
We conduct a sensitivity analysis on the hyperparameter $S$ in the sinusoidal PE enhancement, focusing on its impact on empirical expressivity. We notice that $S$ has no impact on distinguishing basic, regular, and extension graphs, but it displays a stronger influence on distinguishing CFI graphs. In this sensitivity study, we fix the other architectural components and hyperparameters and only vary $S$ along with the size of the first fully-connected layer.
As shown in Table 9, gradually increasing $S$ leads to a better ability to distinguish CFI graphs. This observation matches our motivation for incorporating SPE, as discussed in Sec. 3.4. SPE can effectively enhance the signal differences in graph PE, and thus can make it easier for MLPs to learn how to extract pertinent information. However, the accuracy does not increase monotonically with $S$. Once $S$ is sufficiently large, further increasing it does not necessarily lead to stronger empirical expressivity, and potentially can lead to overfitting and/or demand more training iterations.
Table 9: Sensitivity study on $S$ in Sinusoidal encoding for RRWP on CFI graphs from BREC Wang and Zhang, (2024).
SPE ($S=$) 0 3 6 9 12 15 20
Num. in CFI (100) 3 3 9 17 20 24 22
B.2 Case Study Comparing BN, RMSN and AdaRMSN
We conduct an additional case study to demonstrate the advantages of AdaRMSN in addressing the limitations of token-wise normalization: the ineffectiveness in preserving magnitude information.
We randomly generate a set of points in 2D Euclidean space with magnitudes ranging from 0.5 to 1.5. Given the data points, we train auto-encoders with BN/RMSN/AdaRMSN, respectively,
$\mathbf{y}=\mathtt{FC}\circ\mathtt{Norm}\circ\mathtt{FC}(\mathbf{x})$ (11)
where $\mathtt{FC}$ stands for fully-connected layers (i.e., linear layers) and $\mathtt{Norm}$ denotes the normalization layers. We conduct an overfitting test as a sanity check to measure information loss in the normalization layer via the autoencoders’ ability to recover input data points in predictions. Each auto-encoder is trained independently for 5000 epochs in full-batch mode using the AdamW optimizer. As illustrated in Fig. 5, we present visualizations comparing input data points with predictions generated by autoencoders employing BN, RMSN, and AdaRMSN, respectively. The results demonstrate that RMSN fails to preserve magnitude information, while both BN and AdaRMSN successfully maintain this crucial aspect of the data points. We note that BN relies on cumulative moving averages of mini-batch statistics to estimate population parameters, which can make it sensitive to the choices of batch size.
Figure 5: (Case Study of AdaRMSN) Visualization of Input and Pred data points [(1). Input; (2) Predictions w/ BN; (3) Predictions w/ RMSN; (4) Predictions w/ AdaRMSN]. RMSN is ineffective in preserving magnitude information, whereas both BN and AdaRMSN successfully maintain the crucial magnitude information of the data point
B.3 Sensitivity Study of AdaRMSN w.r.t. Batch Size
To better understand the advantages of AdaRMSN over BatchNorm (BN), we conducted a sensitivity study comparing normalization techniques w.r.t. batch sizes, based on PPGT on the ZINC dataset The study (as shown in Fig. 6) showcases the better stability of AdaRMSN compared to the choice of BN used in many previous GTs Kreuzer et al., (2021); Rampášek et al., (2022); Ma et al., (2023).
Figure 6: Test MAE of PPGT on ZINC v.s. Batch Size (BS). # Training epochs are adjusted per batch-size for the same total update steps: $400*\text{BS}/32$. The first $10\%$ epochs are in the warmup stage. AdaRMSN and RMSN demonstrate better stability and less sensitivity to varying batch sizes compared to BN.
B.4 Runtime and GPU Comsumption
We provide a basic statistics on runtime and gpu comsumption of PPGT, comparing to GRIT (as shown in Table 10).
Table 10: Comparison of peak GPU memory usage and per-epoch training time for GRIT and PPGT. Dataset: Peptides-Structure (15K graphs); Model config.: 5 transformer layers, 96 channels, batch size 32. Hardware: a single Nvidia V100 GPU with 32GB memory, supported by 80 Intel Xeon Gold 6140 CPUs running at 2.30GHz
Model GPU Memory (GB) Training Time (Sec/Epoch)
GRIT 29.16 141.60
PPGT 25.07 100.68
Improv. $\sim$14.03$\sim$28.9%
Appendix C Implementation Details
C.1 Attention in Graph Transformers
In SAN Kreuzer et al., (2021) and GRIT Ma et al., (2023), the attention mechanism is implemented with the sparse operations in PyTorch Geometric due to the complicated attention mechanism. For PPGTs, we provide two versions, one based on the sparse operations and another one based on the dense operation with padding. The latter one is typically faster for larger-scale graphs but consumes more memory.
C.2 Initializations of Parameters
Following most recent plain Transformers, we initialize the weights of linear layers in the backbone as well as the prediction heads with truncated normalization with standard deviation $\sigma=0.02$.
For the stems, we utilize Kaiming uniform initialization He et al., (2015) with $a=0$ for hidden layers in MLPs and $a=1$ for output layers in MLPs or standalone linear layers.
For lookup-table-like embedding layer, nn.Embedding, we utilize the default normal initialization with $\sigma=1$.
C.3 Injection of Degree and Graph Order
For all datasets, we inject the $\log$-degree of each node and the $\log$-graph-order (i.e., the number of nodes in the graph) as additional node attributes. Besides RRWP, we inject the reciprocal of degrees for node $i$ and node $j$ as well as the reciprocal of the graph order to $\mathbf{p}^{\prime}_{ij}$ as an extension.
For superpixel datasets (CIFAR10, MNIST), we also include the location of pixel into graph PE.
C.4 Notes on Reproducibility
Our implementation is built upon PyTorch and PyTorch Geometric. For processing graphs, we utilize the scatter operations from PyTorch Geometric, which are known to be non-deterministic for execution on GPUs. Therefore, even with the same random seed, the experimental results of different trials might vary in an acceptable range. This statement is applicable to most existing models.
We conducted the experiments for ZINC with NVIDIA GeForce RTX 4080 super, experiments for PCQM4Mv2 on NVIDIA H100, and the rest experiments with NVIDIA A100.
Appendix D Additional Related Work
D.1 GRIT’s Attention
The attention mechanism in GRIT Ma et al., (2023) adopts the conditional MLP Perez et al., (2018), involves linear projections, elementwise multiplications, and an uncommon non-linearity in the form of a signed-square-root:
$\hat{\mathbf{e}}_{i,j}$ $=\sigma\Big{(}\rho\left(\left(\mathbf{W}_{\text{Q}}\mathbf{x}_{i}+\mathbf{W}_{\text{K}}\mathbf{x}_{j}\right)\odot\mathbf{W}_{\text{Ew}}\mathbf{e}_{i,j}\right)+\mathbf{W}_{\text{Eb}}\mathbf{e}_{i,j}\Big{)}\in\mathbb{R}^{d^{\prime}},$ (12)
$\alpha_{ij}$ $=\text{Softmax}_{j\in\mathcal{V}}(\mathbf{W}_{\text{A}}\hat{\mathbf{e}}_{i,j})\in\mathbb{R},$
$\hat{\mathbf{x}}_{i}$ $=\sum_{j\in\mathcal{V}}\alpha_{ij}\cdot(\mathbf{W}_{\text{V}}\mathbf{x}_{j}+\mathbf{W}_{\text{Ev}}\hat{\mathbf{e}}_{i,j})\in\mathbb{R}^{d},$
where $\sigma$ is a non-linear activation (ReLU by default); $\mathbf{W}_{\text{Q}},\mathbf{W}_{\text{K}},\mathbf{W}_{\text{Ew}},\mathbf{W}_{\text{Eb}}\in\mathbb{R}^{d^{\prime}\times d}$, $\mathbf{W}_{\text{A}}\in\mathbb{R}^{1\times d^{\prime}}$, $\mathbf{W}_{\text{V}}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{\text{Ev}}\in\mathbb{R}^{d\times d^{\prime}}$ are learnable weight matrices; $\odot$ indicates elementwise multiplication; and $\rho(\mathbf{x}):=(\text{ReLU}(\mathbf{x}))^{1/2}-(\text{ReLU}(-\mathbf{x}))^{1/2}$ is the signed-square-root. Note that $\mathbf{e}_{i,j}$ here corresponds to $\mathbf{p}_{ij}$ in our PPGT but it requires update within the attention mechanism. According to Ma et al., (2023), the signed-square-root $\rho$ is necessary to stabilize the training.
D.2 MPNNs
As the most widely used GNNs, message-passing neural networks (MPNNs) Gilmer et al., (2017); Kipf and Welling, (2017); Hamilton et al., (2017); Veličković et al., (2018) learn graphs, following the 1-WL framework Xu et al., (2019). However, there are several known limitations of MPNNs: (1) over-smoothing Li et al., (2018); Oono and Suzuki, (2020); (2) over-squashing and under-reaching Alon and Yahav, (2020); Topping et al., (2022); and (3) limited expressivity bounded by $1$-WL Xu et al., (2019).
Researchers dedicate relentless efforts to overcome the aforementioned limitations and lead to three research directions: (1). graph Transformers; (2) higher-order GNNs; (3) subgraph GNNs.
D.3 Graph Positional Encoding and Structural Encoding
Attention mechanisms are structure-invariant operators which sense no structural information inherently. Therefore, Transformers strongly rely on positional encoding to capture the structural information Vaswani et al., (2017); Dosovitskiy et al., (2021); Su et al., (2024). Designing graph positional encoding is challenging compared to the counterpart in Euclidean spaces, due to the irregular structure and the symmetry to permutation. Widely used graph positional encoding includes absolute PE: LapPE Dwivedi and Bresson, (2021); Huang et al., (2024); and relative PE: shortest-path distance Ying et al., (2021), resistance-distance Zhang et al., 2023b , and RRWP Ma et al., (2023). Recent works Black et al., (2024); Zhang et al., 2024b study the connections between absolute PE and relative PE.
Besides the aforementioned PE, there exist several structural encoding (SE) approaches that aim to enhance MPNNs, e.g., RWSE Dwivedi et al., 2022b , substructure counting Bouritsas et al., (2022), and homomorphism counting Jin et al., (2024). These SEs can effectively improve the empirical performance and/or theoretical expressivity of MPNNs. Although they are not designed specifically for graph Transformers, integrating them into graph Transformers is usually beneficial.
D.4 Higher-order GNNs
Besides graph Transformers, inspired by the $K$-WL Weisfeiler and Leman, (1968) and $K$-Folklore WL Cai et al., (1992) frameworks, $K$-GNNs Morris et al., (2019); Zhou et al., (2023); Maron et al., (2019) uplift GNNs to higher-order, treating a tuple of $K$ nodes as a token and adapting the color refinement algorithms accordingly. Some other higher-order GNNs Bodnar et al., (2022, 2021) are less closely related to $K$-WL but still perform well. $K$-GNNs can reach theoretical expressivity bounded by $K$-WL, but are typically computationally costly, with $O(N^{K})$ computational complexity.
D.5 Subgraph GNNs
Subgraph GNNs Bevilacqua et al., (2022); Frasca et al., (2022); Zhang et al., 2024a ; Zhang and Li, (2021); Huang et al., (2023) focus on improving the expressivity of GNNs beyond 1-WL. Unlike graph Transformers and higher-order GNNs, subgraph GNNs typically do not change the model architecture. Instead, MPNNs are trained using a novel learning pipeline: a graph is split into multiple subgraphs, subgraph representations are learned independently, and then the subgraph representations are merged to obtain the graph representation. Node-level subgraph GNNs are typically bounded by $3$-WL Frasca et al., (2022), while the expressivity of edge-level subgraph GNNs is still under-explored. Huang et al., (2023) reveal that the expressivity of edge-level GNNs might partially go beyond $3$-WL. Note that subgraph GNNs usually have high memory requirements due to the need to save multiple copies for each input graph.
Although most subgraph GNNs use MPNNs as the base model, the potential integration with stronger graph models, e.g., higher-order GNNs and graph Transformers, is still an open question.
Appendix E Theoretical Analysis
E.1 Theoretical Expressivity of PPGT
[…]
The detailed discussion of GD-WL GNNs and its theoretical expressivity can be found in Sec. 4 and Appx. E.3 of Zhang et al., 2023b , as well as in Sec. 5 of Zhang et al., 2024b . The discussion on RRWP can be found in Sec. 3.1.1 of Ma et al., (2023). The empirical study of realized expressivity (in Sec. 5.1) also matches the theoretical expressivity of PPGT.
Based on the GD-WL analysis framework Zhang et al., 2023b , the proof is straightforward. However, we still provide a simple proof for the completeness of the conclusion.
Proposition E.1.
Powerful Plain Graph Transformers (PPGT) with generalized distance (GD) as graph PE are as powerful as GD-WL, when choosing proper functions $\phi$ and $\theta$ and using a sufficiently large number of heads and layers.
For a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the iterative node color update in GD-WL test is defined as:
$\chi^{\ell}_{\mathcal{G}}(v)=\textit{hash}(\{\!\!\{(d_{\mathcal{G}}(v,u),\chi^{\ell-1}_{\mathcal{G}}(u)):u\in\mathcal{V}\}\!\!\})\,.$ (13)
where $d_{\mathcal{G}}(v,u)$ denotes a distance between nodes $v$ and $u$, and $\chi_{G}^{0}(v)$ is the initial color of $v$. The multiset of final node colors $\{\!\!\{\chi_{G}^{L}(v):v\in\mathcal{V}\}\!\!\}$ at iteration $L$ is hashed to obtain a graph color.
Lemma E.2.
(Lemma 5 of  Xu et al., (2019)) For any countable set $\mathcal{X}$, there exists a function $f:\mathcal{X}\to\mathbb{R}^{n}$ such that $h(\hat{\mathcal{X}}):=\sum_{x\in\hat{\mathcal{X}}}f(x)$ is unique for each multiset $\hat{\mathcal{X}}\in\mathcal{X}$ of bounded size. Moreover, for some function $\phi$, any multiset function $g$ can be decomposed as $g(\hat{\mathcal{X}})=\phi(\sum_{x\in\hat{\mathcal{X}}}f(x))$.
Proof of Proposition E.1.
In this proof, we consider shortest-path distance (SPD) as an example of generalized distance (GD), denoted as $\text{d}_{G}^{\text{SPD}}$, which can be directly extended to other GDs such as the resistance distance (RD) Zhang et al., 2023b and RRWP Ma et al., (2023). Note that the choice of GD determines the practical expressiveness of GD-WL.
We consider all graphs with at most $n$ nodes to distinguish in the isomorphism tests. The total number of possible values of $\text{d}_{G}$ is finite and depends on $n$ (upper bounded by $n^{2}$). We define
$\mathcal{D}_{n}=\{\text{d}_{G}^{\text{SPD}}(u,v):G=(\mathcal{V},\mathcal{E}),|\mathcal{V}|\leqslant n,u,v\in\mathcal{V}\}\,,$ (14)
to denote all possible values of $\text{d}_{G}^{\text{SPD}}(u,v)$ for any graphs with at most $n$ nodes. We note that since $\mathcal{D}_{n}$ is a finite set, its elements can be listed as $\mathcal{D}_{n}=\{d_{G,1},\cdots,d_{G,|\mathcal{D}_{n}|}\}$.
Then the GD-WL aggregation at the $\ell$-th iteration in Eq. (13) can be equivalently rewritten as (See Theorem E.3 in Zhang et al., 2023b ):
$\chi^{\ell}_{G}(v):=\text{hash}\Big{(}\chi_{G}^{\ell,1}(v),\chi_{G}^{\ell,2}(v),\cdots,\chi_{G}^{\ell,|\mathcal{D}_{n}|}(v)\Big{)}\,,$
where $\chi_{G}^{\ell,k}(v):=\{\!\!\{\chi_{G}^{\ell-1}(u):u\in\mathcal{V},\text{d}_{G}(u,v)=d_{G,k}\}\!\!\}\,.$ (15)
In other words, for each node $v$, we can perform a color update by hashing a tuple of color multisets determined by the $\text{d}_{G}$. We construct the $k$-th multiset by injectively aggregating the colors of all nodes $u\in\mathcal{V}$ at a specific distance $d_{G,k}$ from node $v$.
Assuming the color of each node $\chi^{l}_{G}(v)$ is represented as a vector $\mathbf{x}_{v}^{(l)}\in\mathbb{R}^{C}$, by setting the query and key projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K}$ as zero matrices and $\theta$ as a zero function following Zhang et al., 2023b , the attention layer of PPGT with URPE of the $h$-th head (Eq. 5) can be written as
$\hat{\mathbf{x}}^{(l),h}_{v}:=\frac{1}{|\mathcal{V}|}\sum_{u\in\mathcal{V}}(\mathbf{W}_{O}^{h}\mathbf{W}_{V}^{h}\mathbf{x}^{(l)}_{u})\cdot\phi^{h}\big{(}\text{d}_{G}(u,v)\big{)}\,.$ (16)
By defining $\phi(d):=\mathbb{I}(d=d_{G,h})$, where $\mathbb{I}:\mathbb{R}\to\mathbb{R}$ is the indicator function, $d_{G,h}\in\mathcal{D}_{n}$ is a pre-determined condition, we can have
$\hat{\mathbf{x}}^{(l),h}_{v}=$ $\frac{1}{|\mathcal{V}|}\sum_{\text{d}_{G}(u,v)=d_{G,h}}\mathbf{x}^{(l)}_{u}\,.$ (17)
where $\mathbf{W}_{O}^{h}\text{ and }\mathbf{W}_{V}^{h}$ are dropped since they can be absorbed into the following feed-forward networks (FFNs). Note that the constant $\frac{1}{|\mathcal{V}|}$ can be extracted with an additional head and injected back to node representation in the following FFNs.
Then, we can invoke Lemma E.2 to establish that each attention head of PPGT can implement an injective aggregating function for $\{\!\!\{\chi_{G}^{l-1}(u):u\in\mathcal{V},d_{G}(u,v)=d_{G,h}\}\!\!\}$. The summation/concatenation of the output from attention heads is an injective mapping of the tuple of multisets $\left(\chi_{G}^{l,1},\cdots,\chi_{G}^{l,|\mathcal{D}_{n}|}\right)$. When any of the linear mappings has irrational weights, the projection will also be injective. Therefore, with a sufficiently large number of attention heads, the multiset representations $\chi_{G}^{l,k},k\in[|\mathcal{D}_{n}|]$ can be inejectively obtained.
Therefore, with a sufficient number of attention heads and a sufficient number of layers, PPGT is as powerful as GD-WL in distinguishing non-isomorphic graphs, which concludes the proof. ∎
It is worth mentioning that the expressivity upper bound of GD-WL (i.e., 2-FWL/3-WL), given by Theorem 4.5 in Zhang et al., 2023b , is based on the usage of generalized distance (e.g., resistance distance) as graph PE. If we add PE that provides finer structural information, PPGT might surpass the 3-WL expressivity upper bound. This is empirically verified in Section 5.1, where we demonstrate using I2GNN to generate PE for PPGT.
E.2 LN and RMSN are Magnitude Invariant
Proposition E.3.
LN and RMSN are magnitude-invariant, i.e., for an input vector $\mathbf{x}\in\mathbb{R}^{D}$ and a positive scalar $c\in\mathbb{R}^{+}$, $\text{LN}(c\cdot\mathbf{x})=\text{LN}(\mathbf{x})$ and $\text{RMSN}(c\cdot\mathbf{x})=\text{RMSN}(\mathbf{x})$.
Proof of Proposition E.3.
From eq. (1), we immediately have:
$\text{RMSN}(c\cdot\mathbf{x})$ $:=\frac{c\cdot\mathbf{x}}{\frac{1}{\sqrt{D}}\|c\cdot\mathbf{x}\|}\cdot\boldsymbol{\gamma}\,,$
[…]
$=\text{RMSN}(\mathbf{x}),\forall c>0\,,$ (18)
which proves the proposition for RMSN. The proof for LN follows the same steps. ∎
Appendix F Limitations and Discussion
As a plain Transformers, PPGTs still require $O(N^{2})$ computational complexity like other plain Transformers.
This limits the ability of PPGTs to handle very large-size inputs, which are known as large-scale graphs in the context of graph learning.
Note that the large-scale graph data is a different concept from large-scale graphs. We provide a more detailed discussion here.
F.1 Large-scale Graph Datasets v.s. Large-scale-graph Datasets
As foundation models become increasingly popular, researchers have grown interested in the models’ capacity to learn from massive volumes of data, a.k.a., large-scale data.
However, in the domain of graph learning, persistent confusion remains when discussing scaling up graph models for large-scale datasets, particularly concerning two distinct concepts: "large-scale datasets" versus "large-scale-graph datasets."
We provide a preliminary clarification regarding them:
  • Large-scale (graph) datasets (a large number of examples, i.e., graphs): This aligns with the conventional understanding of dataset scale in machine learning (e.g., language and vision), where capacity is the crucial factor for learning from the vast volume of training data. This type of dataset is the usual scenario for training large foundation models.
  • Large-scale-graph dataset (many nodes in a graph): This corresponds to the challenges in long-context learning in language tasks and high-resolution image processing in vision tasks. The challenges mainly lie in the computational efficiency and memory consumption of the models. These datasets are usually small-scale in terms of examples, typically containing only one graph per dataset.
Even though these two research directions are both highly important in solving real-world problems. The improvements on "scalability" and on "capacity" of (graph) models are generally two distinct research directions that are often mutually exclusive in practice.
F.2 Efficiency Techniques for Plain Transformers
As plain Transformers, PPGTs can potentially adopt several efficiency techniques developed for plain Transformers in graph-related and other domains, directly or with modifications.
BigBird Zaheer et al., (2020) and Exphormer Shirzad et al., (2023) propose building sparse attention mechanisms through sparse computational graphs, using random graphs and expander graphs for language tasks and graph learning tasks, respectively. They can be directly adopted by PPGTs.
[…]
Ainslie et al., (2023) propose grouped-query attention (GQA) to save the memory bandwidth. It is a more specific design for decoder-only Transformers (for KV-cache), and might not introduce remarkable benefits to encoder-based Transformers, such as ViTs Dosovitskiy et al., (2021); Touvron et al., 2021a ; Liu et al., (2021), Time-series Transformers Nie et al., (2023); Zhang et al., 2024c and most GTs Kreuzer et al., (2021); Ying et al., (2021); Zhang et al., 2023b ; Ma et al., (2023) including PPGTs.
We mention only a few efficiency techniques for plain Transformers here. Owing to the plain Transformer architecture, PPGTs can potentially adopt efficiency techniques developed in other domains and leverage their training advances.
F.3 Impact Statement
This paper presents work whose goal is to advance the field of Geometric/Graph Deep Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@
 To achieve this, we propose to incorporate three simple, minimal, and easy-to-implement modifications to the plain Transformer architecture to construct our Powerful Plain Graph Transformers (PPGT): (1) simplified $L_{2}$ attention for measuring the magnitude closeness among tokens; (2) adaptive root-mean-square normalization to preserve token magnitude information; and (3) a simple MLP-based stem for graph positional encoding.
-Consistent with its theoretical expressivity, PPGT demonstrates noteworthy realized expressivity on the empirical graph expressivity benchmark, comparing favorably to more complicated competitors such as subgraph GNNs and higher-order GNNs.
-Its outstanding empirical performance across various graph datasets also justifies the practical effectiveness of PPGT.
+Consistent with its theoretical expressivity, PPGT demonstrates noteworthy realized expressivity on the empirical graph expressivity benchmark, comparing favorably to more complicated alternatives such as subgraph GNNs and higher-order GNNs.
+Its empirical performance across various graph datasets also justifies the effectiveness of PPGT.
+This finding underscores the versatility of plain Transformer architectures and highlights their strong potential as a unified backbone for multimodal learning across language, vision, and graph domains.
 1 Introduction
-Transformers have achieved excellent performance across various domains, from language Vaswani et al., (2017); Devlin et al., (2019); Brown et al., (2020) to vision Dosovitskiy et al., (2021); Touvron et al., 2021a ; Touvron et al., 2021b , and are well known for their reduced dependency on inductive bias as well as stronger flexibility and generalizability Dosovitskiy et al., (2021).
-The recent success of transformer-based, multi-modal, large language models (LLMs) OpenAI, (2024); Dubey et al., (2024) has also brought the once unattainable dream of building artificial general intelligence closer to reality.
+Transformers have achieved excellent performance across various domains, from language Vaswani et al.
+(2017); Devlin et al.
+(2019); Brown et al.
+(2020) to vision Dosovitskiy et al.
+(2021); Touvron et al.
+(2021a; b), and are well known for their reduced dependency on inductive bias as well as stronger flexibility and generalizability Dosovitskiy et al.
+(2021).
+The recent success of transformer-based, multi-modal, large language models (LLMs) OpenAI (2024); Dubey et al.
+(2024) has also brought the once-unattainable dream of building artificial general intelligence closer to reality.
 However, these modalities are primarily data on Euclidean spaces, while non-Euclidean domains are still under-explored.
-Non-Euclidean domains like graphs are important for describing physical systems with different symmetry properties, from protein structure prediction Jumper et al., (2021) to complex physics simulation Sanchez-Gonzalez et al., (2018).
+Non-Euclidean domains like graphs are important for describing physical systems with different symmetry properties, from protein structure prediction Jumper et al.
+(2021) to complex physics simulation Sanchez-Gonzalez et al.
+(2018).
 Therefore, many researchers have attempted to migrate Transformer architectures for graph learning for the future integration into multimodal foundation models.
 Nevertheless, unlike other Euclidean domains, the characteristics of graphs make the naive migration challenging.
-Following the suboptimal performance of an early attempt Dwivedi and Bresson, (2021), researchers introduced significant architectural modifications to make Transformers perform well for graph learning.
-These include: implicit/explicit message-passing mechanisms Kreuzer et al., (2021); Rampášek et al., (2022), edge-updating Kim et al., (2022); Hussain et al., (2022); Ma et al., (2023), and sophisticated non-SDP (scaled-dot-product) attention designs Chen et al., (2022); Ma et al., (2023).
+Following the suboptimal performance of an early attempt Dwivedi and Bresson (2021), researchers introduced significant architectural modifications to make Transformers perform well for graph learning.
+These include: implicit/explicit message-passing mechanisms Kreuzer et al.
+(2021); Rampášek et al.
+(2022), edge-updating Kim et al.
+(2022); Hussain et al.
+(2022); Ma et al.
+(2023), and sophisticated non-SDP (scaled-dot-product) attention designs Chen et al.
+(2022); Ma et al.
+(2023).
 These complexities hinder the integration of advances for Transformers arising from other domains and impede progress towards potential future unification of multi-modalities.
@@
 Instead of adding significant architectural modifications, we propose three minimalist but effective modifications to empower plain Transformers with advanced capabilities for learning on graphs (as shown in Fig.
-1(a)).
+1).
 This results in our proposed Powerful Plain Graph Transformers (PPGTs), which have remarkable empirical and theoretical expressivity for distinguishing graph structures and superior empirical performance on real-world graph datasets.
@@
 Learning of Graphs and Encoding of Multisets
-[…]
+Graphs are non-Euclidean geometric spaces with irregular structures and symmetry to permutation (i.e., invariant/equivariant).
+A key factor for learning graphs is the ability to distinguish the distinct structures of the input graphs, which is typically referred to as the expressivity of the graph model Xu et al.
+(2019); Zhang et al.
+(2023b).
+Currently, most graph neural networks (GNNs) are developed based on the framework of the Weisfeiler-Leman (WL) Isomorphism test Weisfeiler and Leman (1968) – a color-refinement algorithm based on multisets $\{\!\{\cdot\}\!\}$ encoding.
+For example, message-passing networks (MPNNs) based on 1-WL Xu et al.
+(2019), distinguish graphs by encoding the neighborhood of each node as a multiset Xu et al.
+(2019).
+To go beyond 1-WL, researchers have extended to $K$-WL GNNs Morris et al.
+(2019), $K$-Folklore-WL (FWL) GNNs Feng et al.
+(2023), and Generalized-distance-WL (GD-WL) GNNs Zhang et al.
+(2023b), which are also based on such multiset encoding schemes, albeit with different multiset objects.
+It is worth mentioning that, as discussed in Xu et al.
+(2019); Zhang et al.
+(2023b), the cardinalities of multisets are crucial for distinguishing multisets, e.g., $\{\!\{a,b\}\!\}$ versus $\{\!\{a,a,b,b\}\!\}$.
+Generally, the cardinality is encoded into the token representation as its magnitude, e.g., encoding $\{\!\{a,b\}\!\}$ and $\{\!\{a,a,b,b\}\!\}$ as $\mathbf{x}$ and $c\cdot\mathbf{x}$ respectively, for $c\neq 1\in\mathbb{R}^{+}$.
+Thus, the loss of token magnitude information weakens the ability to distinguish multisets and, consequently, graph structures.
 Graph Transformers – Learning Graphs with Pseudo-coordinates
 The philosophy of Graph Transformers (GTs) is to learn graph representations using positional encodings (PEs) rather than operating directly on the original input graphs.
-Zhang et al., 2023b provide a theoretical framework (GD-WL) that demonstrates that GTs’ stronger expressivity stems from finer-grained information beyond 1-WL encoded in graph PEs.
-Zhang et al., 2024b demonstrate that GD-WL not only applies to distance-like or relative graph PEs but also enables analysis of absolute PEs (e.g., Laplacian PE) by transforming them into their relative counterparts.
-Concurrently, Ma et al., (2024) interpret graph PE as pseudo-coordinates in graph spaces that mimic manifolds, demonstrating that it can extend beyond graph distance and is not restricted to the Transformer architecture (e.g., convolutions on pseudo-coordinates).
-We follow the same philosophy and show that with a simple stem on pseudo-coordinates, even plain Transformers can achieve outstanding expressive power in graph learning.
+Zhang et al.
+(2023b) provide a theoretical framework (GD-WL) that demonstrates that GTs’ stronger expressivity stems from finer-grained information beyond 1-WL encoded in graph PEs.
+Zhang et al.
+(2024b) demonstrate that GD-WL not only applies to distance-like or relative graph PEs but also enables analysis of absolute PEs (e.g., Laplacian PE) by transforming them into their relative counterparts.
+Concurrently, Ma et al.
+(2024) interpret graph PE as pseudo-coordinates in graph spaces that mimic manifolds, demonstrating that it can extend beyond graph distance and is not restricted to the Transformer architecture (e.g., convolutions on pseudo-coordinates).
+We follow the same philosophy and show that with a simple stem on pseudo-coordinates, even plain Transformers can achieve strong expressive power in graph learning.
 2.2 Limitations in Plain Transformer Architectures
 The Loss of Magnitude Information in Token-wise Normalization Layer
-LayerNorm (LN Ba et al., (2016)) and Root-Mean-Square-Norm (RMSN Zhang and Sennrich, (2019)) are two widely used token-wise normalization techniques in Transformer-based models that effectively control token magnitudes:
-$\text{LN}(\mathbf{x})=\frac{\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}\|}\odot\boldsymbol{\gamma}+\boldsymbol{\beta}\,,\quad$ $\text{RMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\odot\boldsymbol{\gamma}\,.$ (1)
-[…]
+LayerNorm (LN Ba et al.
+(2016)) and Root-Mean-Square-Norm (RMSN Zhang and Sennrich (2019)) are two widely used token-wise normalization techniques in Transformer-based models that effectively control token magnitudes:
+$\text{LN}(\mathbf{x})=\frac{\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}\|}\odot\bm{\gamma}+\bm{\beta}\,$$\quad$ $\text{RMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\odot\bm{\gamma}\,.$ (1)
+Here $\mathbf{x}\in\mathbb{R}^{D}$ are token vectors; $\|\cdot\|\in[0,\infty)$ is the $L_{2}$-norm (i.e., magnitude) of a vector; and $\bm{\gamma},\bm{\beta}\in\mathbb{R}^{D}$ are parameters of a learnable affine transform.
+They retract token representations onto a hypersphere, a property essential for dot-product attention mechanisms, echoing the notion of retraction as used in computational physics.
+However, both LN and RMSN are strictly invariant to changes in input magnitude (see Proposition E.3 and the case study in Appx.
+B.2), which can result in the loss of valuable token magnitude information
 Pitfalls of Scaled Dot-product Attention
-With a good balance of capacity and efficiency, Scaled dot-product (SDP) attention has become the most foundational attention mechanism in modern Transformers Vaswani et al., (2017).
+With a good balance of capacity and efficiency, Scaled dot-product (SDP) attention has become the most common attention mechanism in modern Transformers Vaswani et al.
+(2017).
 It has been widely explored in previous works and well optimized in deep learning libraries.
@@
 Cos attention disregards the magnitude information.
-$L_{2}$ attention strikes a balance among SDP and Cos attention to attend to $k_{1}$, which has the lowest $L_{2}$ distance to the query $q$.
+$L_{2}$ attention strikes a balance between SDP and Cos attention to attend to $k_{1}$, which has the lowest $L_{2}$ distance to the query $q$.
 Three drawbacks of SDP attention necessitate the additional control of token magnitudes:
   1. 1.
-    Softmax saturation: large token magnitudes (i.e., $\|\mathbf{q}_{i}\|$, $\|\mathbf{k}_{j}\|$) lead to large pre-softmax-logit values $\hat{\alpha}_{ij}$ and consequently extremely small gradients in backpropagation Vaswani et al., (2017).
+    Softmax saturation: large token magnitudes (i.e., $\|\mathbf{q}_{i}\|$, $\|\mathbf{k}_{j}\|$) lead to large pre-softmax-logit values $\hat{\alpha}_{ij}$ and consequently extremely small gradients in backpropagation Vaswani et al.
+(2017).
   2. 2.
@@
 To mitigate the first and third drawbacks, existing plain Transformers heavily rely on token-wise normalization to regulate the token magnitudes.
-Removing LN/RMSN and/or using BatchNorm(BN) Ioffe and Szegedy, (2015) instead can induce training instability and divergence Touvron et al., 2021b ; Yao et al., (2021).
+Removing LN/RMSN and/or using BatchNorm(BN) Ioffe and Szegedy (2015) instead can induce training instability and divergence Touvron et al.
+(2021b); Yao et al.
+(2021).
 3 Method
 As established in Section 2, plain Transformers exhibit suboptimal performance in graph learning, due to the architectural limitations that hinder preserving and modeling token magnitude information — a critical aspect of capturing graph structures.
-Consequently, previous GTs attempt to address it by adopting BN and overly sophisticated attention designs (e.g., MPNNs and/or MLP-based attention) Kreuzer et al., (2021); Rampášek et al., (2022); Ma et al., (2023).
+Consequently, previous GTs attempt to address it by adopting BN and overly sophisticated attention designs (e.g., MPNNs and/or MLP-based attention) Kreuzer et al.
+(2021); Rampášek et al.
+(2022); Ma et al.
+(2023).
 They have thus strayed away from plain Transformers, impeding the transfer of previously explored training advances and the potential unification of other foundation models.
@@
 Furthermore, we introduce an extra enhancement in the PE stem to boost the information extraction of PE.
-These enhancements enable plain Transformers to achieve outstanding expressive power for graph learning.
+These enhancements enable plain Transformers to achieve stronger expressive power for graph learning.
 3.1 AdaRMSN
@@
 We desire a normalization layer that can not only control the token magnitudes, but is also capable of preserving the magnitude information when necessary.
-Inspired by adaptive normalization layers Dumoulin et al., (2017); de Vries et al., (2017); Peebles and Xie, (2023), we propose adaptive RMSN (AdaRMSN)
-$\text{AdaRMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\cdot\gamma^{\prime}(\mathbf{x})\,,\quad\text{where }\gamma^{\prime}(\mathbf{x}):=\frac{1}{\sqrt{D}}\|\boldsymbol{\alpha}\cdot\mathbf{x}+\boldsymbol{\beta}\|\,.$ (3)
-The parameter $\boldsymbol{\beta}\in\mathbb{R}^{D}$ is initialized as $\mathbf{1}$ and $\boldsymbol{\alpha}\in\mathbb{R}^{D}$ as $\mathbf{0}$, leading to $\gamma^{\prime}(\mathbf{x})=1$.
-AdaRMSN behaves the same as regular RMSN at the initial stage of training, but is capable of recovering the identity transformation with $\boldsymbol{\beta}=\mathbf{0}$ and $\boldsymbol{\alpha}=\mathbf{1}$ when necessary.
+Inspired by adaptive normalization layers Dumoulin et al. (2017); de Vries et al. (2017); Peebles and Xie (2023), we propose adaptive RMSN (AdaRMSN)
+$\text{AdaRMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\cdot\gamma^{\prime}(\mathbf{x})\,,\quad\text{where }\gamma^{\prime}(\mathbf{x}):=\frac{1}{\sqrt{D}}\|\bm{\alpha}\cdot\mathbf{x}+\bm{\beta}\|\,.$ (3)
+The parameter $\bm{\beta}\in\mathbb{R}^{D}$ is initialized as $\mathbf{1}$ and $\bm{\alpha}\in\mathbb{R}^{D}$ as $\mathbf{0}$, leading to $\gamma^{\prime}(\mathbf{x})=1$.
+AdaRMSN behaves the same as regular RMSN at the initial stage of training, but is capable of recovering the identity transformation with $\bm{\beta}=\mathbf{0}$ and $\bm{\alpha}=\mathbf{1}$ when necessary.
 3.2 Simplified $L_{2}$ Attention Mechanisms
 From Dot-product to Euclidean Distance
-To empower SDP attention to sense both angle- and magnitude-information among query and key tokens, we revisit $L_{2}$ attention Kim et al., (2021), which is based on Euclidean distance and is capable of measuring token closeness by balancing both angles and magnitudes, in contrast to SDP attention and cosine-similarity attention (as shown in Fig.
+To empower SDP attention to sense both angle- and magnitude-information among query and key tokens, we revisit $L_{2}$ attention Kim et al.
+(2021), which is based on Euclidean distance and is capable of measuring token closeness by balancing both angles and magnitudes, in contrast to SDP attention and cosine-similarity attention (as shown in Fig.
 2).
@@
 $=\mathtt{Softmax}_{j}(\frac{1}{\sqrt{D}}\mathbf{q}_{i}^{\intercal}\mathbf{k}_{j}-\frac{1}{2\sqrt{D}}\mathbf{k}_{j}^{\intercal}\mathbf{k}_{j})\,.$ (4)
-This can be easily supported by existing SDP attention implementations in most deep learning libraries, e.g., PyTorch Paszke et al., (2019) 111e.g., the bias term can be directly formulated as a float attention mask in SDP-attention in PyTorch..
+This can be easily supported by existing SDP attention implementations in most deep learning libraries, e.g., PyTorch Paszke et al.
+(2019) 111e.g., the bias term can be directly formulated as a float attention mask in SDP-attention in PyTorch..
 […]
@@
 where $\phi:\mathbb{R}^{D}\to\mathbb{R}$ and $\theta:\mathbb{R}^{D}\to\mathbb{R}$ are linear transforms.
-The $\phi(\mathbf{\mathbf{p}_{ij}})$ is an optional term purely based on the relative position to guarantee the universality of attention with relative PE (URPE) Luo et al., (2022).
-This term demands slight customization of the existing attention implementation, but we retain it since it is beneficial for learning objects with complicated structures Luo et al., (2022); Zhang et al., 2023b .
-Note that, the attention with PE in relative-form is general, since it is widely employed in many existing Transformers, from language Shaw et al., (2018); Raffel et al., (2020); Press et al., (2022) to vision Dosovitskiy et al., (2021); Liu et al., (2021, 2022).
-Many absolute PEs are de facto explictly/implictly transformed into the relative-form in attention mechanisms Su et al., (2024); Huang et al., (2024); Zhang et al., 2024b .
+The $\phi(\mathbf{\mathbf{p}_{ij}})$ is an optional term purely based on the relative position to guarantee the universality of attention with relative PE (URPE) Luo et al.
+(2022).
+This term demands slight customization of the existing attention implementation, but we retain it since it is beneficial for learning objects with complicated structures Luo et al.
+(2022); Zhang et al.
+(2023b).
+Notably, this form can be viewed as Continuous Kernel Graph Convolution Ma et al.
+(2024) with a dynamic density function.
+Note that, the attention with PE in relative-form is general, since it is widely employed in many existing Transformers, from language Shaw et al.
+(2018); Raffel et al.
+(2020); Press et al.
+(2022) to vision Dosovitskiy et al.
+(2021); Liu et al.
+(2021; 2022).
+Many absolute PEs are de facto explicitly/implicitly transformed into the relative-form in attention mechanisms Su et al.
+(2024); Huang et al.
+(2024); Zhang et al.
+(2024b).
 3.3 Powerful Plain Graph Transformers
 In this section, we provide an example of constructing plain Transformers for learning graphs with our proposed techniques, termed Powerful Plain Graph Transformers (PPGT).
-We follow the plain Vision Transformers architecture Dosovitskiy et al., (2021); Touvron et al., 2021a , — stems, backbone and prediction head (as shown in Fig. 1(a)(a) and Fig. 1(a)(d)).
+We follow the plain Vision Transformers architecture Dosovitskiy et al. (2021); Touvron et al. (2021a), — stems, backbone and prediction head (as shown in Fig. 1(a) and Fig. 1(d)).
 Graph Positional Encoding
-In this work, we utilize relative random walk probabilities (RRWP) Ma et al., (2023) as our demonstrating example of graph PE, considering its simplicity and effectiveness. RRWP is defined as
-$\mathbf{p}^{\prime}_{ij}=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\dots,\mathbf{M}^{K-1}]_{[i,j]}\in\mathbb{R}^{K}\,,$ (6)
-where $\mathbf{X}_{[i,j]}$ stands for the $i,j$th element/slice of a tensor $\mathbf{X}$; $\mathbf{M}:=\mathbf{D}^{-1}\mathbf{A}$ is the random walk matrix given the adjacency matrix $\mathbf{A}$ of the graph; and $\mathbf{I}\in\mathbb{R}^{N\times N}$ denotes the identity matrix.
+In this work, we utilize relative random walk probabilities (RRWP) Ma et al. (2023) as our demonstrating example of graph PE, considering its simplicity and effectiveness. RRWP is defined as
+$\mathbf{p}^{\prime}_{ij}=[\mathbf{I},{\textnormal{w}},{\textnormal{w}}^{2},\dots,{\textnormal{w}}^{K-1}]_{[i,j]}\in\mathbb{R}^{K}\,$, (6)
+where $\mathbf{X}_{[i,j]}$ stands for the $i,j$th element/slice of a tensor $\mathbf{X}$; ${\textnormal{w}}:=\mathbf{D}^{-1}\mathbf{A}$ is the random walk matrix given the adjacency matrix $\mathbf{A}$ of the graph; and $\mathbf{I}\in\mathbb{R}^{N\times N}$ denotes the identity matrix.
 Most other graph PEs are applicable in our framework with minor modifications to the PE stem.
 Transformer Backbone
-Following the latest plain Transformers, we utilize the pre-norm Xiong et al., (2020) architecture, for $l=1,\cdots,L$,
+Following the latest plain Transformers, we utilize the pre-norm Xiong et al. (2020) architecture, for $l=1,\cdots,L$,
 $\hat{\mathbf{X}}^{l}=\mathbf{X}^{l-1}+\mathtt{MSA}(\mathtt{Norm}(\mathbf{X}^{l-1}),\mathbf{P}),\quad\mathbf{X}^{l}=\mathtt{FFN}(\mathbf{X}^{l}):=\hat{\mathbf{X}}^{l}+\mathtt{MLP}(\mathtt{Norm}(\hat{\mathbf{X}}^{l})),\quad\mathbf{Y}=\mathtt{Norm}(\mathbf{X}^{L})$ (7)
@@
 Stem
-Before the Transformer backbone, we use small networks, usually referred to as stems, to process positional encoding $\mathbf{p}^{\prime}_{ij}$ and merge node/edge attributes $\mathbf{x}^{\prime}_{i}$/$\mathbf{e}^{\prime}_{ij}$.
+Before the Transformer backbone, we use small networks, usually referred to as stems, to process positional encoding $\mathbf{p}^{\prime}_{ij}$ and merge node/edge attributes $\mathbf{x}^{\prime}_{i}$$\mathbf{e}^{\prime}_{ij}$.
 We consider a simple stem design for node and PE, respectively:
@@
 where $\mathtt{FC}$ stands for the fully-connected layer (e.g., linear projection); $\circ$ stands for function composition; $\mathbf{x}^{\prime}_{i}$ is treated as zero if there are no node attributes; $\mathbf{e}^{\prime}_{ij}$ is set to zero if there are no edge attributes or if $(i,j)$ is not an observed edge.
-Driven by the analysis of RRWP in Ma et al., (2023), we introduce additional FFNs and a final normalization layer in the PE stem to better extract the structural information, mimicking the pre-norm architecture of the Transformer backbone.
+Driven by the analysis of RRWP in Ma et al.
+(2023), we introduce additional FFNs and a final normalization layer in the PE stem to better extract the structural information, mimicking the pre-norm architecture of the Transformer backbone.
 Prediction Head
-We use a task-specific MLP prediction head, following GRIT Ma et al., (2023) and GraphGPS Rampášek et al., (2022): sum/mean-pooling followed by a MLP for graph level tasks; and a MLP shared by all nodes for node level tasks.
-3.4 Sinusoidal PE Enhancement (SPE)
+Unless otherwise specified, we employ a task-specific MLP prediction head, following the designs of GRIT Ma et al.
+(2023) and GraphGPS Rampášek et al.
+(2022) – for graph-level tasks, we apply sum or mean pooling followed by an MLP; for node-level tasks, we use an MLP shared across all nodes.
+For OGBN-ArXiv, we utilize a class-attention prediction head inspired by CaiT Touvron et al.
+(2021b), which is better suited for the graph-sampling strategy.
+3.4 Anti-Spectral-Bias with Sinusoidal PE Enhancement (SPE)
 Figure 3: Illustration of two node-pairs (a) $(i,j)$ and $(i,k)$ of a graph, and (b) absolute difference of RRWPs and sinusoidally-encoded RRWPs for those two node-pairs.
-As discussed in Zhang et al., 2023b ; Zhang et al., 2024b , the finer-grained the information provided through graph positional encodings (PEs), the stronger distinguishability of graph structure GD-GNNs can achieve.
-However, due to the spectral bias of neural networks Rahaman et al., (2019), MLPs prioritize learning the low-frequency modes and thus lose detailed information stored in PE.
-To mitigate this issue, motivated by NeRF Mildenhall et al., (2020), we add an extra sinusoidal encoding on top of the RRWP and process it with a simple MLP.
+As discussed in Zhang et al.
+(2023b; 2024b), as the information provided through graph positional encodings (PEs) becomes more fine-grained, GD-GNNs can achieve better distinguishability of graph structure.
+However, due to the spectral bias of neural networks Rahaman et al.
+(2019), MLPs prioritize learning the low-frequency modes and thus lose detailed information stored in PEs.
+To mitigate this issue, motivated by NeRF Mildenhall et al.
+(2020), we add an extra sinusoidal encoding on top of the RRWP and process it with a simple MLP.
 The sinusoidal encoding is applied to each channel of $\mathbf{p}^{\prime}_{ij}$ in an elementwise fashion:
@@
 The PE stem (Eq.
-8) becomes
+(8)) becomes
 $\mathbf{p}^{0}_{ij}=\mathtt{Norm}\circ\mathtt{FFN}\circ\cdots\circ\mathtt{FFN}(\mathtt{FC}(\mathbf{e}^{\prime}_{ij})+\mathtt{MLP}(\mathtt{SPE}(\mathbf{p}^{\prime}_{ij})))$ (10)
@@
 3.5 Theoretical Expressivity of PPGT
-The theoretical expressivity of PPGT can be analyzed with the GD-WL framework Zhang et al., 2023b . Since it is not proposed in our work, we provide further discussion of it in Appx. E.1.
+The theoretical expressivity of PPGT can be analyzed within the GD-WL framework Zhang et al.
+(2023b).
+When equipped with regular generalized distances (e.g., RRWP, resistance distance Zhang et al.
+(2023b)) as graph PE, PPGT attains expressivity strictly beyond $1$-WL and is upper bounded by $3$-WL, matching the theoretical limits of a broad class of GD-WL algorithms.
+We provide the formal proof in Appx.
+E.1.
 4 Relationship with Previous Work
-Graph Transformers Motivated by the success of Transformers in other domains, researchers have strived to migrate Transformers to graph learning to address the limitations of MPNNs.
-Although the naive migration of Transformer to graph learning did not work well Dwivedi and Bresson, (2021), several recent works have achieved considerable success when constructing graph Transformers, from theoretical expressivity analysis Zhang et al., 2023b ; Zhang et al., 2024b to outstanding empirical performance Ying et al., (2021); Rampášek et al., (2022); Ma et al., (2023).
+Graph Transformers.
+Transformers, especially plain Transformers—architectures close to the vanilla Transformer Vaswani et al.
+(2017) with scaled dot-product (SDP) attention and feed-forward networks (FFNs)—have achieved outstanding performance across a wide range of domains, from language Devlin et al.
+(2019); Raffel et al.
+(2020); OpenAI (2024); Dubey et al.
+(2024) to vision Dosovitskiy et al.
+(2021); Touvron et al.
+(2021b); Caron et al.
+(2021); Oquab et al.
+(2024).
+Motivated by the success of Transformers in other domains, researchers have strived to migrate Transformers to graph learning to address the limitations of MPNNs.
+Although the naive migration of Transformers to graph learning did not work well Dwivedi and Bresson (2021), several recent works have achieved considerable success when constructing graph Transformers, from theoretical expressivity analysis Zhang et al.
+(2023b; 2024b) to impressive empirical performance Ying et al.
+(2021); Rampášek et al.
+(2022); Ma et al.
+(2023).
 However, during the development of these graph Transformers, there has been a gradual but substantial deviation from the plain Transformers widely used in other domains.
-For example, SAN Kreuzer et al., (2021) introduces dual-attention mechanisms with local and global aggregations; K-Subgraph SAT Chen et al., (2022) introduce MPNNs into attention mechanisms to compute attention scores; GraphGPS Rampášek et al., (2022) heavily relies on the MPNNs within its hybrid Transformer architecture; EGT Hussain et al., (2022) introduces gated mechanisms and edge-updates inside the attention; GRIT Ma et al., (2023) incorporates a complicated conditional MLP-based attention mechanism (shown in Appendix.
+For example, SAN Kreuzer et al.
+(2021) introduces dual-attention mechanisms with local and global aggregations; K-Subgraph SAT Chen et al.
+(2022) introduces MPNNs into attention mechanisms to compute attention scores; GraphGPS Rampášek et al.
+(2022) heavily relies on the MPNNs within its hybrid Transformer architecture; EGT Hussain et al.
+(2022) introduces gated mechanisms and edge-updates inside the attention; GRIT Ma et al.
+(2023) incorporates a complicated conditional MLP-based attention mechanism (shown in Appendix.
 D.1).
 These deviations prevent the easy adoption of Transformer training advances and obscure the potential unification of cross-modality foundation models.
-Among these graph Transformers, the graphormer-series Ying et al., (2021); Luo et al., (2022); Zhang et al., 2023b retain architectures that are closest to plain Transformers; unfortunately, the result is a substantial gap in empirical performance and empirical expressivity to the best-performing graph Transformers Rampášek et al., (2022); Ma et al., (2023).
-In this work, we analyze the limitations of plain Transformers for learning on graphs and propose simple but effective designs to allow plain transformers to reach outstanding empirical performance as well as remarkable expressivity.
-Other Attention Mechanisms Besides SDP attention, there are other attention variants in use.
-Bahdanau et al., (2015) introduced the earliest content-based attention mechanism for recurrent neural networks (RNNs) based on an MLP, which is more computational and memory costly.
-Swin-Transformer-V2 Liu et al., (2022) uses cosine-similarity to compute attention scores for better stability, but neglects the magnitude information.
-Kim et al., (2021) propose the use of the negative of the square $L_{2}$-distance, with tied query-key projection weights, for maintaining Lipchitz continuity of Transformers.
+Among these graph Transformers, the graphormer-series Ying et al.
+(2021); Luo et al.
+(2022); Zhang et al.
+(2023b) retain architectures that are closest to plain Transformers; unfortunately, the result is a substantial gap in empirical performance and empirical expressivity compared to the best-performing graph Transformers Rampášek et al.
+(2022); Ma et al.
+(2023).
+Most graph Transformers treat a node as a token.
+TokenGT Kim et al.
+(2022) views both nodes and edges as tokens and processes them using plain or sparse Transformers.
+However, despite adopting a plain Transformer architecture, TokenGT also falls considerably behind recent graph Transformers on the expressivity-demanding PCQM4Mv2 dataset.
+Developing powerful graph Transformers based on a plain Transformer architecture is particularly attractive, as the associated hardware stacks and software libraries have already been extensively optimized.
+Therefore, we investigate the fundamental limitations of plain Transformers on graph-structured data and introduce several simple yet effective architectural enhancements that preserve the core plain Transformer design, enabling competitive empirical performance without requiring significant architectural modifications.
+Other Attention Mechanisms.
+Besides SDP attention, there are other attention variants in use.
+Bahdanau et al.
+(2015) introduced the earliest content-based attention mechanism for recurrent neural networks (RNNs) based on an MLP, which is more computationally and memory costly.
+Swin-Transformer-V2 Liu et al.
+(2022) uses cosine-similarity to compute attention scores for better stability, but neglects the magnitude information.
+Kim et al.
+(2021) propose the use of the negative of the square $L_{2}$-distance, with tied query-key projection weights, for maintaining Lipchitz continuity of Transformers.
 Our attention mechanism, although based on $L_{2}$ attention, is further simplified and adjusted in order to maintain alignment with SDP attention.
-Additional related work We discuss additional related work, including MPNNs, graph positional/structural encoding, higher-order GNNs, and subgraph GNNs in Appx. D.
+Continuous Kernel Graph Convolution.
+Ma et al.
+(2024) introduce graph convolution operators with continuous kernels defined over pseudo-coordinates of graphs, termed CKGConv.
+These operators offer better flexibility in capturing high-frequency information compared to attention-based mechanisms.
+PPGT with URPE enhancement can also be interpreted as a generalization of CKGConv.
+In Eq.
+(5), the $\phi(\mathbf{p}_{ij})$ term serves as the convolution kernel, analogous to that in CKGConv, while the $\mathtt{Softmax}_{j}\big(\frac{\mathbf{q}_{i}^{\intercal}\mathbf{k}_{j}}{\sqrt{D}}-\frac{\mathbf{k}_{j}^{\intercal}\mathbf{k}_{j}}{2\sqrt{D}}+\theta(\mathbf{p}_{ij})\big)$ component can be regarded as a dynamic density function conditioned on the token representations, which is assumed to be uniform in CKGConv.
+Universality of Transformers The Universal Approximation Theorem is fundamental for understanding the theoretical representational capacity upper bound of neural networks.
+While Hornik et al.
+(1989) established that MLPs are universal approximators for functions $f:\mathbb{R}^{N}\to\mathbb{R}^{M}$ on compact sets, this conclusion does not automatically extend to other domains.
+Yun et al.
+(2020) demonstrated that Transformers with absolute Positional Encodings (PE)—which sufficiently distinguish token order—are universal approximators for sequence-to-sequence functions.
+Conversely, Luo et al.
+(2022) showed that Transformers with relative PE lack this universality.
+However, they show that it can be restored through the universality enhancement.
+Regarding graph domains, Kreuzer et al.
+(2021) state that Transformers with absolute PE can approximate any function $f$ for the graph isomorphism problem.
+Crucially, however, this requires that the absolute PE can uniquely identify nodes in each graph, which is infeasible due to the highly symmetric structure of graphs.
+Graph Transformers with full attention and appropriate positional encoding (PE) are typically considered at most as expressive as the 3-WL test for graph isomorphism Zhang et al.
+(2023b).
+In contrast, those using linear attention have expressivity equivalent to MPNNs with virtual nodes Cai et al.
+(2023).
+Additional related work. We discuss additional related work, including MPNNs, graph positional/structural encoding, higher-order GNNs, and subgraph GNNs in Appx. D.
 5 Experimental Results
 5.1 Empirical Expressivity on Graph Isomorphism
-To better understand the expressivity of PPGT, we evaluate our model on the BREC benchmark Wang and Zhang, (2024), a comprehensive dataset for measuring the empirical expressive power of GNNs w.r.t.
+To better understand the expressivity of PPGT, we evaluate our model on the BREC benchmark Wang and Zhang (2024), a comprehensive dataset for measuring the empirical expressive power of GNNs w.r.t.
 graph isomorphism, with graph-pairs from 1-WL to 4-WL-indistinguishable.
@@
 Mismatch between expressivity and real-world benchmark performance: The stronger theoretical/empirical expressivity is not completely reflected in real-world benchmark performance.
-For example, subgraph GNNs and/or K-WL GNNs with stronger expressivity (e.g., SSWL+, I2GNN, N2GNN) reach inferior performance compared to GRIT and PPGT on the ZINC benchmark.
+For example, subgraph GNNs and/or K-WL GNNs with stronger expressivity (e.g., SSWL+, I2GNN, N2GNN) demonstrate inferior performance compared to GRIT and PPGT on the ZINC benchmark.
 [4].
-Going beyond GD-WL?: As previously discussed, the design of graph PE, a.k.a., pseudo-coordiantes, can be extended beyond the distance/affinity of graphs.
-With this in mind, we conduct a exploratory demo called I2GNN+PPGT, which uses I2GNN to generate additional positional encodings for PPGT.
+Going beyond GD-WL?: As previously discussed, the design of graph PE, a.k.a., pseudo-coordinates, can be extended beyond the distance/affinity of graphs.
+With this in mind, we conduct an exploratory demo called I2GNN+PPGT, which uses I2GNN to generate additional positional encodings for PPGT.
 The empirical expressivity is further improved to 76%, outperforming the standalone I2GNN and PPGT, surpassing 3-WL and reaching the top performance among the methods compared.
 This demonstration hints that plain Graph Transformers can potentially surpass GD-WL and achieve greater expressive power purely through enhanced positional encoding designs.
+This result is noteworthy, as PPGT achieves superior empirical performance compared to many subgraph GNNs and higher-order GNNs that possess greater theoretical expressivity.
+The expressivity bottleneck of our PPGT model does not stem from its architecture, but rather from the design of positional encodings (PE).
+To fully realize the expressive potential of plain Graph Transformers, it is crucial to develop expressive and generalizable PE schemes for graphs that are also compatible with permutation symmetry.
 […]
@@
 […]
-[…]
+I2GNN $\cdot$ $\not\sqsupseteq$3-WL 60 100 100 21 281 70.2% 0.083
 K-WL GNNs PPGN 3-WL 60 50 100 23 233 58.2% -
@@
 […]
-I2GNN+PPGT GD+WL? 60 120 100 24 304 76% -
+I2GNN+PPGT GD++-WL 60 120 100 24 304 76% -
 5.2 Benchmarking PPGT on Real-world Benchmarks
-The specifications, details, and references for the baseline methods are provided in Appx. A.5.
+The specifications, details, and references for the baseline methods are provided in Appx. A.6.
 Benchmarking GNNs
-We conduct a general evaluation of our proposed PPGT on five datasets from Benchmarking GNNs (Dwivedi et al., 2022a, ): ZINC, MNIST, CIFAR10, PATTERN, and CLUSTER, and summarize the results in Table 2.
+We conduct a general evaluation of our proposed PPGT on five datasets from Benchmarking GNNs (Dwivedi et al., 2022a): ZINC, MNIST, CIFAR10, PATTERN, and CLUSTER, and summarize the results in Table 2.
 We observe that our model obtains the best mean performance for all five datasets, outperforming various MPNNs, non-MPNN GNNs, and existing graph Transformers.
 These results showcase the effectiveness of PPGT for general graph learning with a plain Transformer architecture.
-Table 2: Test performance in five benchmarks from Benchmarking GNNs (baselines please see Appx.
-A.5 for details and references).
+Table 2: Test performance on five benchmarks from Benchmarking GNNs (baselines please see Appx.
+A.6 for details and references).
 Shown is the mean $\pm$ s.d.
@@
 GatedGCN $0.282\pm 0.015$ $97.340\pm 0.143$ $67.312\pm 0.311$ $85.568\pm 0.088$ $73.840\pm 0.326$
-PNA $0.188\pm 0.004$ $97.94\pm 0.12$ $70.35\pm 0.63$- --
+PNA $0.188\pm 0.004$ $97.94\pm 0.12$ $70.35\pm 0.63$ - -
 Non-MPNN Graph Neural Networks
-CRaW1 $0.085\pm 0.004$ ${97.944}\pm{0.050}$ $69.013\pm 0.259$- --
+CRaW1 $0.085\pm 0.004$ ${97.944}\pm{0.050}$ $69.013\pm 0.259$ - -
 GIN-AK+ ${0.080}\pm{0.001}$ - $72.19\pm 0.13$ $86.850\pm 0.057$ -
@@
 Graph Transformers
-SAN $0.139\pm 0.006$- -- $86.581\pm 0.037$ $76.691\pm 0.65$
-K-Subgraph SAT $0.094\pm 0.008$- -- ${86.848\pm 0.037}$ $77.856\pm 0.104$
+SAN $0.139\pm 0.006$ - - $86.581\pm 0.037$ $76.691\pm 0.65$
+K-Subgraph SAT $0.094\pm 0.008$ - - $86.848\pm 0.037$} $77.856\pm 0.104$
 EGT $0.108\pm 0.009$ $\mathbf{98.173\pm 0.087}$ $68.702\pm 0.409$ $86.821\pm 0.020$ $\mathbf{79.232\pm 0.348}$
-Graphormer-GD $0.081\pm 0.009$- -- -- --
+Graphormer-GD $0.081\pm 0.009$ - - - -
 GPS $0.070\pm 0.004$ $98.051\pm 0.126$} $72.298\pm 0.356$} $86.685\pm 0.059$ $78.016\pm 0.180$}
-GMLP-Mixer $0.077\pm 0.003$- -- -- --
+GMLP-Mixer $0.077\pm 0.003$ - - - -
 GRIT $\mathbf{0.059\pm 0.002}$ $98.108\pm 0.111$ $\mathbf{76.468\pm 0.881}$ $\mathbf{87.196\pm 0.076}$ $\mathbf{80.026\pm 0.277}$
-PGT (Ours) $\mathbf{0.0566\pm 0.002}$ $\mathbf{98.614\pm 0.096}$ $\mathbf{78.560\pm 0.700}$ $\mathbf{89.752\pm 0.030}$ $\mathbf{80.027\pm 0.114}$
+PPGT $\mathbf{0.0566\pm 0.002}$ $\mathbf{98.614\pm 0.096}$ $\mathbf{78.560\pm 0.700}$ $\mathbf{89.752\pm 0.030}$ $\mathbf{80.027\pm 0.114}$
 Long Range Graph Benchmarks
-We present experimental results on two Long-range Graph Benchmark (LRGB) Dwivedi et al., 2022c datasets, Peptides-Function and Peptides-Structure in Table 5.2.
-PPGT achieves the lowest MAE on Peptides-Structure and remains in the top three models on Peptides-Function.
+We present experimental results on three Long-range Graph Benchmark (LRGB) Dwivedi et al.
+(2022c) datasets – Peptides-Function, Peptides-Structure and PASCALVOC-SP in Table 3.
+PPGT achieves the lowest MAE on Peptides-Structure, top F1 on PascalVoc-SP and remains in the top three models on Peptides-Function.
+We adopt the updated experimental setup from LRGB as described in Tönshoff et al.
+(2024), and report the corresponding results.
 Moreover, the difference in AP for Peptides-Function dataset between the best performing GRIT and our PPGT is not statistically significant at the 5% level for a one-tailed t-test.
-These results demonstrate PPGT’s outstanding capability of learning long-range dependency structures.
-PCQM4Mv2 Large-scale Graph Regression Benchmark
-Further, we conduct an experiment on the PCQM4Mv2 large-scale graph regression benchmark [^2] Clarification on large-scale graph-datasets and large-scale-graph datasets can be found in Appx.
-F.1.
-of 3.7M graphs Hu et al., (2021), which is currently one of the largest molecular datasets (Table 5.2).
-We follow the experimental setting of Rampášek et al., (2022) — excluding 3D information as attributes, and treating the validation set of PCQM4Mv2 as the test set, instead of the original private Test-dev set.
-The result of a single random seed run is reported due to the size of the dataset, following previous works.
-We followed the hyperparameter configuration of GRIT Ma et al., (2023) and did not conduct hyperparameter search due to computational limits.
-However, our model can still reach an outstanding performance, comparable to GRIT and GraphGPS, with similar parameter budget.
-Table 3: Test performance on two benchmarks from Long-range Graph Benchmarks (LRGB). (mean $\pm$ s.d. of 4 runs). Highlighted are the top first, second, and third results.
-Model Peptides-func Peptides-struct
-[…]
-GCN $0.5930\pm 0.0023$ $0.3496\pm 0.0013$
-GINE $0.5498\pm 0.0079$ $0.3547\pm 0.0045$
-GatedGCN+RWSE $0.6069\pm 0.0035$ $0.3357\pm 0.0006$
-CKGCN ${0.6952\pm 0.0068}$ ${0.2477\pm 0.0018}$
-Transformer+LapPE $0.6326\pm 0.0126$ $0.2529\pm 0.0016$
-SAN+LapPE $0.6384\pm 0.0121$ $0.2683\pm 0.0043$
-SAN+RWSE ${0.6439\pm 0.0075}$ $0.2545\pm 0.0012$
-GPS $0.6535\pm 0.0041$ $0.2500\pm 0.0012$
-GMLP-Mixer $\mathbf{0.6970\pm 0.0080}$ $\mathbf{0.2475\pm 0.0015}$
-GRIT $\mathbf{0.6988\pm 0.0082}$ $\mathbf{0.2460\pm 0.0012}$
-PPGT (Ours) $\mathbf{0.6961\pm 0.0062}$ $\mathbf{0.2450\pm 0.0017}$
+These results demonstrate PPGT’s capability of learning long-range dependency structures.
+5.3 Benchmarking PPGT on Large-scale Graph Benchmark and Large-scale-graph Benchmark
+Scaling behavior with respect to both model size and dataset size is a critical consideration.
+Accordingly, in this section, we evaluate the applicability of PPGT to large-scale data settings without imposing a strict parameter budget.
+We consider two distinct types of large-scale data settings of graphs: large-scale graph benchmarks and large-scale-graph benchmarks.
+Although prior work often conflates these two scenarios, the distinction is important.
+Specifically, large-scale (graph) datasets comprise many graph instances, requiring scalability with respect to dataset size.
+In contrast, large-scale graph datasets consist of a single graph with an extremely large number of nodes, necessitating scalability with respect to input size (analogous to Gigapixel image processing).
+A more detailed discussion is provided in Appendix F.1.
+Large-scale Graph Benchmark: PCQM4Mv2
+To further assess the scalability of PPGT to large-scale data, we evaluate its performance on the PCQM4Mv2 large-scale graph regression benchmark, which consists of 3.7 million graphs Hu et al.
+(2021).
+This dataset is among the largest-scale graph benchmarks to date (see Table 5.3).
+Following the experimental protocol of Rampášek et al.
+(2022), we exclude the 3D information from the model attributes and use the PCQM4Mv2 validation set in place of the original private Test-dev set for evaluation.
+By omitting 3D information, we aim to more accurately measure each model’s intrinsic graph learning capability, ensuring that results do not reflect reliance on 3D coordinates.
+The result of a single random seed run is reported due to the size of the dataset, following previous works. Our model performs well, comparably to GRIT and GraphGPS, with a similar parameter budget.
+Table 3: Performance comparison on Long-Range Graph Benchmark (LRGB)Peptides and PascalVoc-SP datasets. (mean $\pm$ s.d. of 4 runs). Highlighted are the top first, second, and third results.
+Method Peptides-Func Peptides-Struct PascalVoc-SP
+AP $\uparrow$ MAE $\downarrow$ F1 $\uparrow$
+GCN $0.6860\pm 0.0050$ $\mathbf{0.2460\pm 0.0007}$ $0.2078\pm 0.0031$
+GINE $0.6621\pm 0.0067$ $\mathbf{0.2473\pm 0.0017}$ $0.2718\pm 0.0054$
+GatedGCN $0.6765\pm 0.0047$ $0.2477\pm 0.0009$ $0.3880\pm 0.0040$
+DRew $\mathbf{0.7150\pm 0.0044}$ $0.2536\pm 0.0015$ $0.3314\pm 0.0024$
+Exphormer $0.6527\pm 0.0043$ $0.2481\pm 0.0007$ $\mathbf{0.3960\pm 0.0027}$
+GPS $0.6534\pm 0.0090$ $0.2509\pm 0.0010$ $\mathbf{0.4440\pm 0.0065}$
+GRIT $\mathbf{0.6988\pm 0.0082}$ $\mathbf{0.2460\pm 0.0012}$ -
+PPGT $\mathbf{0.6961\pm 0.0062}$ $\mathbf{0.2450\pm 0.0017}$ $\mathbf{0.4641\pm 0.0033}$
 Table 4: Test performance on PCQM4Mv2 dataset. Shown is the result of a single run due to the computation constraint. Highlighted are the top first, second, and third results.
@@
 PPGT (Ours) $\mathbf{0.0856}$ 17.6M
+Large-scale-graph Benchmark: OGBN-ArXiv
+Unlike large-scale graph benchmarks, which prioritize the model’s capacity to learn from many graph instances, a large-scale-graph benchmark typically emphasizes memory efficiency to handle the substantial size of an individual graph.
+Even though PPGT is inherently not designed for processing large-size inputs, we also benchmark our approach on a large-scale-graph dataset: OGBN-ArXiv Hu et al.
+(2020) [^2] No feature enhancement techniques, such as UniMP Shi et al.
+(2021) or GIANT Chien et al.
+(2022), are applied..
+To handle large graphs, we adopt an additional graph sampling strategy.
+Specifically, we convert node-level tasks into graph-level tasks by extracting a local inducted subgraph around each target node using breadth-first search (BFS) node sampling, referred to as Node2Subgraph conversion.
+Similar techniques exist Zeng et al.
+(2020); Sun et al.
+(2023), but we do not explore them here, as they are beyond the focus of this work.
+Even in efficiency-oriented graph Transformers, graph sampling techniques are widely adopted to handle ultra–large-scale input graphs that cannot be processed directly.
+We compare PPGT with several state-of-the-art efficiency-oriented graph Transformers, including NodeFormer Wu et al.
+(2022), Exphormer Shirzad et al.
+(2023), and SGFormer Wu et al.
+(2023).
+These models aim for full node coverage on large graphs but often compromise expressivity, as linear-attention graph Transformers exhibit the equivalent theoretical expressivity as MPNNs with virtual nodes Cai et al.
+(2023).
+PPGT maintains high expressivity through Node2Subgraph sampling, even at the cost of reduced node coverage, and achieves performance comparable to these SOTA efficiency-oriented graph Transformers.
+These results suggest that both model expressivity and node coverage are essential for learning on large graphs, and achieving a balance between the two is more important than overemphasizing either.
+Table 5: Testing results (Accuracy) on OGBN-ArXiv (mean $\pm$ s.d.
+of 4 runs).
+Highlighted are the top first, second, and third results.
+Baseline results from  (Wu et al., 2023; Shirzad et al., 2023; Li et al., 2022).
+(*DeeperGCN does not provide s.d.)
+Method GCN
+GCN-
+NSampler
+DeeperGCN SIGN NodeFormer SGFormer Exphormer PPGT
+OGBN-ArXiv $71.74\pm 0.29$ $68.50\pm 0.23$ $71.90$ $70.28\pm 0.25$ $59.90\pm 0.42$ $72.63\pm 0.13$ $72.44\pm 0.28$ $72.46\pm 0.15$
 5.4 Ablation Study on Proposed Designs
+[…]
 Figure 4: Ablation Study on ZINC. MLPA: Conditional MLP Attention; DegS: degree scaler; ARMSN: AdaRMSNorm; URP: Universal RPE; SPE: Sinusoidal PE enhancement.
-[…]
-Further Study
-We present further experimental studies and results in Appx. B.
+5.5 Further Study
+In Appx.
+B, we present additional experimental analyses, including: (B.1) a sensitivity study on the number of bases $S$ in SPE, demonstrating the impact of SPE on empirical expressivity; (B.2) a case study of different normalization layers, showing that RMSN discards magnitude information whereas AdaRMSN preserves it; (B.3) a batch-size sensitivity analysis of AdaRMSN, demonstrating its robustness to batch size, in contrast to BN; and (B.4) a comparison of runtime and GPU memory consumption between PPGT and GRIT.
 6 Conclusion
@@
 In this work, we demonstrate that plain Transformers can be powerful graph learners via the proposed minimal, easy-to-adapt modifications.
-Our Powerful Plain Graph Transformers (PPGT) not only achieve competitive expressivity, but also demonstrate outstanding empirical performance on real-world graph benchmarks while maintaining the simplicity of plain Transformer architectures.
+Our Powerful Plain Graph Transformers (PPGTs) not only achieve competitive expressivity, but also demonstrate strong empirical performance on real-world graph benchmarks while maintaining the simplicity of plain Transformer architectures.
 Further exploration of empirical expressivity also unveils a potential direction for improving plain GTs.
 We consider this work an important first step toward reducing dissimilarities between GTs and Transformers in other domains.
-This direction of investigation is crucial for building general multi-modal foundation models that integrate learning capabilities on graphs and potentially other irregular non-Euclidean geometric spaces.
+This potentially eases the design of general multi-modal foundation models that integrate learning capabilities on graphs and potentially other irregular non-Euclidean geometric spaces.
+Beyond graph domains, the insights from our work may also benefit Transformer architectures in other domains, broadening their applicability and impact.
 Limitations:
-This work demonstrates that plain Transformers can be powerful graph learners.
-However, like most plain Transformers, PPGTs still require the $O(N^{2})$ computational complexity and might encounter issues on scaling to large-size inputs.
-A more detailed discussion of limitations and possible avenues to address them is provided in Appx.
+This work demonstrates that plain Transformers can serve as powerful graph learners without external graph models (e.g., MPNNs).
+However, like most vanilla Transformer architectures, PPGTs incur an $\mathcal{O}(N^{2})$ computational complexity, which limits their direct applicability to large-size input graphs.
+To mitigate this issue, graph sampling techniques can be employed to handle large inputs more efficiently.
+A more detailed discussion of these limitations, along with potential directions for addressing them, is provided in Appx.
 F.
 Appendix A Experimental Details
-A.1 Bechmarking GNNs and Long Range Graph Benchmarks
+A.1 Benchmarking GNNs and Long Range Graph Benchmarks
 Description of Datasets
-Table 5 provides a summary of the statistics and characteristics of datasets used in this paper.
-The first five datasets are from Dwivedi et al., 2022a , and the last two are from Dwivedi et al., 2022c .
-Readers are referred to Rampášek et al., (2022) for more details of the datasets.
-Table 5: Overview of the graph learning datasets involved in this work Dwivedi et al., 2022a ; Dwivedi et al., 2022c ; Irwin et al., (2012); Hu et al., (2021).
+Table 6 provides a summary of the statistics and characteristics of the datasets used in this paper.
+The first five datasets are from Dwivedi et al.
+(2022a), and the last two are from Dwivedi et al.
+(2022c).
+Readers are referred to Rampášek et al.
+(2022) for more details of the datasets.
+Table 6: Overview of the graph learning datasets involved in this work Dwivedi et al. (2022a; c); Irwin et al. (2012); Hu et al. (2021).
 Dataset # Graphs Avg. # nodes Avg. # edges Directed Prediction level Prediction task Metric
@@
 Peptides-struct 15,535 150.9 307.3 No Graph 11-task regression Mean Abs. Error
+PascalVoc-SP 11,355 479.4 2,710.5 No Inductive Node 21-class classif. Marco F1
 PCQM4Mv2 3,746,620 14.1 14.6 No Graph regression Mean Abs. Error
-Dataset splits, random seed and parameter budgets
-We conduct the experiments on the standard train/validation/test splits of the evaluated benchmarks, following previous works Rampášek et al., (2022); Ma et al., (2023).
+OGBN-ArXiv 1 169,343 1,116,243 Yes Transductive Node 40-class classif. Accuracy
+Dataset splits, random seed, and parameter budgets
+We conduct the experiments on the standard train/validation/test splits of the evaluated benchmarks, following previous works Rampášek et al.
+(2022); Ma et al.
+(2023).
 For each dataset, we execute 4 trials with different random seeds (0, 1, 2, 3) and report the mean performance and standard deviation.
@@
 Due to the limited time and computational resources, we did not perform an exhaustive search on the hyperparameters.
-We start with the hyperparameter setting of GRIT Ma et al., (2023) and perform minimal search to satisfy the commonly used parameter budgets.
+We start with the hyperparameter setting of GRIT Ma et al.
+(2023) and perform minimal search to satisfy the commonly used parameter budgets.
 The hyperparameters are presented in Table 7 and Table 8.
 In the tables, S.D.
-stands for stochastic depth Huang et al., (2016), a.k.a., drop-path, in which we treat a graph as one example in stochastic depth.
+stands for stochastic depth Huang et al.
+(2016), a.k.a., drop-path, in which we treat a graph as one example in stochastic depth.
 For Peptides-func, we find that there exists a mismatch between the cross-entropy loss and the metric, average precision, due to the highly imbalanced label distribution.
 Empirically, we observe that adding a BN in the prediction head post-graph-pooling effectively mitigates the problem.
-Table 6: Hyperparameters for five datasets from Benchmarking GNNs Dwivedi et al., 2022a
+Table 7: Hyperparameters for five datasets from Benchmarking GNNs Dwivedi et al. (2022a)
 Hyperparameter ZINC MNIST CIFAR10 PATTERN CLUSTER
@@
 MLP-dim 128 64 64 128 128
+Edge-dim 64 32 32 64 64
 # FFN 2 1 2 2 2
 FFN expansion 2 2 2 2 2
+PE-dim 24 10 10 32 32
 # bases 3 3 3 3 3
@@
 # layers 3 2 2 2 2
-PE-dim 24 10 10 32 32
 Training
@@
 # parameters 487K 102K 108K 497K 496K
-Table 7: Hyperparameters for two datasets from the Long-range Graph Benchmark Dwivedi et al., 2022c
-Hyperparameter Peptides-func Peptides-struct
+Table 8: Hyperparameters for two datasets from the Long-range Graph Benchmark Dwivedi et al. (2022c)
+Hyperparameter Peptides-func Peptides-struct PascalVoc-SP
 PE Stem
-MLP-dim 128 128
-# FFN 1 1
-FFN expansion 2 2
-# bases 3 3
+MLP-dim 128 128 128
+Edge-dim 64 64 64
+# FFN 1 1 1
+FFN expansion 2 2 2
+# bases 3 3 3
+PE-dim 32 24 16
 Backbone
-# blocks 5 5
-Dim 96 96
-FFN expansion 2 2
-# attn. heads 16 8
-S.D. 0.1 0.1
-Attn. dropout 0.2 0.1
+# blocks 5 5 12
+Dim 96 96 64
+FFN expansion 2 2 2
+# attn. heads 16 8 8
+S.D. 0.1 0.1 0.2
+Attn. dropout 0.2 0.1 0.3
 Pred. Head
-Graph Pooling sum mean
-Norm BN -
-# layers 3 2
-PE-dim 32 24
+Graph Pooling sum mean -
+Norm BN - -
+# layers 3 2 2
 Training
-Batch size 32 32
-Learn. rate 7e-4 7e-4
-# epochs 400 250
-# warmup 10 10
-Weight decay 1e-5 1e-5
-# Parameters 509K 488K
+Batch size 32 32 16
+Learn. rate 7e-4 7e-4 1e-3
+# epochs 400 250 200
+# warmup 10 10 10
+Weight decay 1e-5 1e-5 1e-2
+# Parameters 509K 488K 492K
 A.2 BREC: Empirical Expressivity Benchmark
-BREC Wang and Zhang, (2024) is an empirical expressivity datasets, consisting of four major categories of graphs: Basic, Regular, Extension, and CFI.
+BREC Wang and Zhang (2024) is an empirical expressivity dataset, consisting of four major categories of graphs: Basic, Regular, Extension, and CFI.
 Basic graphs include $60$ pairs of simple 1-WL-indistinguishable graphs.
 Regular graphs include $140$ pairs of regular graphs from four types of subcategories: $50$ pairs of simple regular graphs, $50$ pairs of strongly regular graphs, $20$ pairs of 4-vertex condition graphs and $20$ pairs of distance regular graphs.
-Extension graphs include $100$ pairs of special graphs that arise when comparing four kinds of GNN extensions Papp and Wattenhofer, (2022).
-CFI graphs include $100$ pairs of graphs generated by CFI methods Cai et al., (1992).
+Extension graphs include $100$ pairs of special graphs that arise when comparing four kinds of GNN extensions Papp and Wattenhofer (2022).
+CFI graphs include $100$ pairs of graphs generated by CFI methods Cai et al.
+(1992).
 All pairs of graphs are 1-WL-indistinguishable.
-We follow the standard training pipelines from BREC: pairwise contrastive training and evaluate process. No specific parameter budget required on BREC dataset.
+We follow the standard training pipelines from BREC: pairwise contrastive training and evaluating process. No specific parameter budget required on the BREC dataset.
 Hyperparameter
@@
 We employ a 6-layer GIN with BN on each subgraph independently to get its representation.
-The subgraph representations are fed to PPGT as the additional edge-attributes.
+The subgraph representations are fed to PPGT as additional edge-attributes.
 The I2-GNN and PPGT are trained end-to-end together.
-Table 8: Hyperparameters for BREC Wang and Zhang, (2024)
+Table 9: Hyperparameters for PCQM4Mv2 Hu et al. (2021) and OGBN-ArXiv Hu et al. (2020).
+Hyperparameter PCQM4Mv2 OGBN-ArXiv
+PE Stem
+MLP-dim 512 192
+Edge-dim 128 96
+# FFN 8 2
+FFN expansion 2 2
+# bases 10 3
+PE-dim 24 5 (dual direction)
+Backbone
+# blocks 16 10+2(class attention)
+Dim 512 192
+FFN expansion 2 2
+# attn. heads 16 8
+S.D. 0.2 0.5
+Attn. dropout 0.2 0.5
+Node2Subgraph
+Node Masking - 0.5
+Max Size of Graph - 100
+Pred. Head
+Graph Pooling sum class attention
+Norm - -
+# layers 3 1
+Training
+Batch size 2048 128
+Learn. rate 1e-3 2e-3
+# epochs 500 300
+# warmup 50 10
+Weight decay 1e-3 5e-2
+# parameters 17.6M 3.74M
+A.3 PCQM4Mv2 from OGB Large-Scale Challenge
+PCQM4Mv2 is a large-scale quantum chemistry graph dataset benchmark, containing over 3.7M graphs, proposed from OGB Large-Scale Challenge (OGB-LSC) Hu et al.
+(2021).
+The statistics of the dataset can be found in Table 6.
+A.4 OGBN-ArXiv from OGB Benchmark
+OGBN-ArXiv is a large-scale graph benchmark featuring a transductive node classification task from Open Graph Benchmark Hu et al.
+(2020).
+Unlike the other graph datasets used in this work, OGBN-ArXiv comprises a single large-scale graph containing over 169,000 nodes.
+This benchmark emphasizes the graph models’ scalability to large input graphs.
+The statistics of the dataset can be found in Table 6.
+For PPGT, we convert the transductive node-classification task on OGBN-ArXiv into a graph-classification setting via the Node2Subgraph transformation.
+In this framework, we employ the class-attention mechanism Touvron et al.
+(2021b) as the output head.
+As OGBN-ArXiv is a directed graph, we compute dual-direction RRWP features, i.e., random walks along both in-edge and out-edge directions.
+Table 10: Hyperparameters for BREC Wang and Zhang (2024)
 Hyperparameter BREC
@@
 MLP-dim 192
+Edge-dim 96
 # FFN 4
@@
 # parameters 874K
-A.3 PCQM4Mv2 from OGB Large-Scale Challenge
-PCQM4Mv2 is a large-scale quantum chemistry graph dataset benchmark, containing over 3.7M graphs, proposed from OGB Large-Scale Challenge (OGB-LSC) Hu et al., (2021).
-The statistics of the dataset can be found in Table 5.
-A.4 Optimizer and Learning Rate Scheduler
-Following most plain Transformers in other domains, we use AdamW Loshchilov and Hutter, (2019) as the optimizer and the Cosine Annealing Learning Rate scheduler Loshchilov and Hutter, (2017) with linear warm up.
-A.5 Baselines Information
-For benchmarks from BenchmarkingGNN Dwivedi et al., 2022a .
-  • Tyical Message-passing Networks(MPNNs): GCN Kipf and Welling, (2017), GIN Xu et al., (2019), GAT Veličković et al., (2018), GatedGCN Bresson and Laurent, (2018), PNA Corso et al., (2020);
-  • GNNs going beyond MPNNs: CRaW1 Tönshoff et al., (2023), GIN-AK+ Zhao et al., (2022), DGN Beani et al., (2021), CKGCN Ma et al., (2024),
-  • Graph Transformers SAN Kreuzer et al., (2021), K-Subgraph SAT Chen et al., (2022), EGT Hussain et al., (2022), Graphormer-GD Zhang et al., 2023b , GPS Rampášek et al., (2022), GMLP-Mixer He et al., (2023), GRIT Ma et al., (2023).
-For BREC Wang and Zhang, (2024):
-  • Subgraph GNNs: SUN Frasca et al., (2022), SSWL+ Zhang et al., 2023a , I2-GNN Huang et al., (2023)
-  • K-WL/K-FWL GNNs: PPGN Maron et al., (2019), 2-DRFWL(2) Zhou et al., (2023), 3-DRFWL(2) Zhou et al., (2023)
-  • K-FWL+SubgraphGNNs: N2GNN Feng et al., (2023)
-  • GD-WL GNNs Graphormer Ying et al., (2021), EPNN Zhang et al., 2024b , CKGConv Ma et al., (2024), GRIT Ma et al., (2023).
-For OGB-LSC Hu et al., (2021):
-MPNNs (GCN Kipf and Welling, (2017), GIN Xu et al., (2019) with/without virtual nodes) as well as several Graph Transformers (GRPE Park et al., (2022), Graphormer Ying et al., (2021), TokenGT Kim et al., (2022) and GraphGPS Rampášek et al., (2022)).
+A.5 Optimizer and Learning Rate Scheduler
+Following most plain Transformers in other domains, we use AdamW Loshchilov and Hutter (2019) as the optimizer and the Cosine Annealing Learning Rate scheduler Loshchilov and Hutter (2017) with linear warm up.
+A.6 Baselines Information
+For benchmarks from BenchmarkingGNN Dwivedi et al. (2022a).
+  • Tyical Message-passing Networks(MPNNs): GCN Kipf and Welling (2017), GIN Xu et al. (2019), GAT Veličković et al. (2018), GatedGCN Bresson and Laurent (2018), PNA Corso et al. (2020);
+  • GNNs going beyond MPNNs: CRaW1 Tönshoff et al. (2023), GIN-AK+ Zhao et al. (2022), DGN Beani et al. (2021), CKGCN Ma et al. (2024),
+  • Graph Transformers SAN Kreuzer et al.
+(2021), K-Subgraph SAT Chen et al.
+(2022), EGT Hussain et al.
+(2022), Graphormer-GD Zhang et al.
+(2023b), GPS Rampášek et al.
+(2022), GMLP-Mixer He et al.
+(2023), GRIT Ma et al.
+(2023).
+For BREC Wang and Zhang (2024):
+  • Subgraph GNNs: SUN Frasca et al. (2022), SSWL+ Zhang et al. (2023a), I2-GNN Huang et al. (2023)
+  • K-WL/K-FWL GNNs: PPGN Maron et al. (2019), 2-DRFWL(2) Zhou et al. (2023), 3-DRFWL(2) Zhou et al. (2023)
+  • K-FWL+SubgraphGNNs: N2GNN Feng et al. (2023)
+  • GD-WL GNNs Graphormer Ying et al. (2021), EPNN Zhang et al. (2024b), CKGConv Ma et al. (2024), GRIT Ma et al. (2023).
+For OGB-LSC Hu et al. (2021):
+MPNNs (GCN Kipf and Welling (2017), GIN Xu et al.
+(2019) with/without virtual nodes) as well as several Graph Transformers (GRPE Park et al.
+(2022), Graphormer Ying et al.
+(2021), TokenGT Kim et al.
+(2022) and GraphGPS Rampášek et al.
+(2022)).
 Appendix B Additional Study
@@
 Once $S$ is sufficiently large, further increasing it does not necessarily lead to stronger empirical expressivity, and potentially can lead to overfitting and/or demand more training iterations.
-Table 9: Sensitivity study on $S$ in Sinusoidal encoding for RRWP on CFI graphs from BREC Wang and Zhang, (2024).
+Table 11: Sensitivity study on $S$ in Sinusoidal encoding for RRWP on CFI graphs from BREC Wang and Zhang (2024).
 SPE ($S=$) 0 3 6 9 12 15 20
-Num. in CFI (100) 3 3 9 17 20 24 22
+# Correct in CFI (100) 3 3 9 17 20 24 22
 B.2 Case Study Comparing BN, RMSN and AdaRMSN
@@
 We note that BN relies on cumulative moving averages of mini-batch statistics to estimate population parameters, which can make it sensitive to the choices of batch size.
-Figure 5: (Case Study of AdaRMSN) Visualization of Input and Pred data points [(1).
-Input; (2) Predictions w/ BN; (3) Predictions w/ RMSN; (4) Predictions w/ AdaRMSN].
+Figure 5: (Case Study of AdaRMSN) Visualization of Input and Pred data points [(1) Input; (2) Predictions w/ BN; (3) Predictions w/ RMSN; (4) Predictions w/ AdaRMSN].
 RMSN is ineffective in preserving magnitude information, whereas both BN and AdaRMSN successfully maintain the crucial magnitude information of the data point
@@
 batch sizes, based on PPGT on the ZINC dataset The study (as shown in Fig.
-6) showcases the better stability of AdaRMSN compared to the choice of BN used in many previous GTs Kreuzer et al., (2021); Rampášek et al., (2022); Ma et al., (2023).
+6) showcases the better stability of AdaRMSN compared to the choice of BN used in many previous GTs Kreuzer et al.
+(2021); Rampášek et al.
+(2022); Ma et al.
+(2023).
 Figure 6: Test MAE of PPGT on ZINC v.s.
@@
 AdaRMSN and RMSN demonstrate better stability and less sensitivity to varying batch sizes compared to BN.
-B.4 Runtime and GPU Comsumption
-We provide a basic statistics on runtime and gpu comsumption of PPGT, comparing to GRIT (as shown in Table 10).
+B.4 Runtime and GPU Consumption
+We report basic statistics on the runtime and GPU memory consumption of PPGT, in comparison with GRIT (see Table 12).
+The results suggest that, even without efficiency-driven techniques, adopting a plain Transformer architecture—rather than more complex graph-specific Transformer designs used in prior work—can yield improved runtime and reduced GPU memory consumption.
 Table 12: Comparison of peak GPU memory usage and per-epoch training time for GRIT and PPGT.
@@
 PPGT 25.07 100.68
-Improv. $\sim$14.03$\sim$28.9%
+Improv. $\sim$14.03% $\sim$28.9%
 Appendix C Implementation Details
 C.1 Attention in Graph Transformers
-In SAN Kreuzer et al., (2021) and GRIT Ma et al., (2023), the attention mechanism is implemented with the sparse operations in PyTorch Geometric due to the complicated attention mechanism.
-For PPGTs, we provide two versions, one based on the sparse operations and another one based on the dense operation with padding.
+In SAN Kreuzer et al.
+(2021) and GRIT Ma et al.
+(2023), the attention mechanism is implemented with the sparse operations in PyTorch Geometric due to the complicated attention mechanism.
+For PPGTs, we provide two versions, one based on the sparse operations and another one based on the dense operations with padding.
 The latter one is typically faster for larger-scale graphs but consumes more memory.
 C.2 Initializations of Parameters
-Following most recent plain Transformers, we initialize the weights of linear layers in the backbone as well as the prediction heads with truncated normalization with standard deviation $\sigma=0.02$.
-For the stems, we utilize Kaiming uniform initialization He et al., (2015) with $a=0$ for hidden layers in MLPs and $a=1$ for output layers in MLPs or standalone linear layers.
+Following the most recent plain Transformers, we initialize the weights of linear layers in the backbone as well as the prediction heads with truncated normalization with standard deviation $\sigma=0.02$.
+For the stems, we utilize Kaiming uniform initialization He et al. (2015) with $a=0$ for hidden layers in MLPs and $a=1$ for output layers in MLPs or standalone linear layers.
 For lookup-table-like embedding layer, nn.Embedding, we utilize the default normal initialization with $\sigma=1$.
@@
 Besides RRWP, we inject the reciprocal of degrees for node $i$ and node $j$ as well as the reciprocal of the graph order to $\mathbf{p}^{\prime}_{ij}$ as an extension.
-For superpixel datasets (CIFAR10, MNIST), we also include the location of pixel into graph PE.
+For superpixel datasets (CIFAR10, MNIST), we also include the location of pixel into the graph PE.
 C.4 Notes on Reproducibility
@@
 D.1 GRIT’s Attention
-The attention mechanism in GRIT Ma et al., (2023) adopts the conditional MLP Perez et al., (2018), involves linear projections, elementwise multiplications, and an uncommon non-linearity in the form of a signed-square-root:
+The attention mechanism in GRIT Ma et al.
+(2023) adopts the conditional MLP Perez et al.
+(2018), involves linear projections, elementwise multiplications, and an uncommon non-linearity in the form of a signed-square-root:
 $\hat{\mathbf{e}}_{i,j}$ $=\sigma\Big(\rho\left(\left(\mathbf{W}_{\text{Q}}\mathbf{x}_{i}+\mathbf{W}_{\text{K}}\mathbf{x}_{j}\right)\odot\mathbf{W}_{\text{Ew}}\mathbf{e}_{i,j}\right)+\mathbf{W}_{\text{Eb}}\mathbf{e}_{i,j}\Big)\in\mathbb{R}^{d^{\prime}},$ (12)
@@
 where $\sigma$ is a non-linear activation (ReLU by default); $\mathbf{W}_{\text{Q}},\mathbf{W}_{\text{K}},\mathbf{W}_{\text{Ew}},\mathbf{W}_{\text{Eb}}\in\mathbb{R}^{d^{\prime}\times d}$, $\mathbf{W}_{\text{A}}\in\mathbb{R}^{1\times d^{\prime}}$, $\mathbf{W}_{\text{V}}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{\text{Ev}}\in\mathbb{R}^{d\times d^{\prime}}$ are learnable weight matrices; $\odot$ indicates elementwise multiplication; and $\rho(\mathbf{x}):=(\text{ReLU}(\mathbf{x}))^{1/2}-(\text{ReLU}(-\mathbf{x}))^{1/2}$ is the signed-square-root.
-Note that $\mathbf{e}_{i,j}$ here corresponds to $\mathbf{p}_{ij}$ in our PPGT but it requires update within the attention mechanism.
-According to Ma et al., (2023), the signed-square-root $\rho$ is necessary to stabilize the training.
+Note that $\mathbf{e}_{i,j}$ here corresponds to $\mathbf{p}_{ij}$ in our PPGT, but it requires updating within the attention mechanism.
+According to Ma et al.
+(2023), the signed-square-root $\rho$ is necessary to stabilize the training.
 D.2 MPNNs
-As the most widely used GNNs, message-passing neural networks (MPNNs) Gilmer et al., (2017); Kipf and Welling, (2017); Hamilton et al., (2017); Veličković et al., (2018) learn graphs, following the 1-WL framework Xu et al., (2019).
-However, there are several known limitations of MPNNs: (1) over-smoothing Li et al., (2018); Oono and Suzuki, (2020); (2) over-squashing and under-reaching Alon and Yahav, (2020); Topping et al., (2022); and (3) limited expressivity bounded by $1$-WL Xu et al., (2019).
+As the most widely used GNNs, message-passing neural networks (MPNNs) Gilmer et al.
+(2017); Kipf and Welling (2017); Hamilton et al.
+(2017); Veličković et al.
+(2018) learn graphs, following the 1-WL framework Xu et al.
+(2019).
+However, there are several known limitations of MPNNs: (1) over-smoothing Li et al.
+(2018); Oono and Suzuki (2020); (2) over-squashing and under-reaching Alon and Yahav (2020); Topping et al.
+(2022); and (3) limited expressivity bounded by $1$-WL Xu et al.
+(2019).
 Researchers dedicate relentless efforts to overcome the aforementioned limitations and lead to three research directions: (1). graph Transformers; (2) higher-order GNNs; (3) subgraph GNNs.
@@
 Attention mechanisms are structure-invariant operators which sense no structural information inherently.
-Therefore, Transformers strongly rely on positional encoding to capture the structural information Vaswani et al., (2017); Dosovitskiy et al., (2021); Su et al., (2024).
+Therefore, Transformers strongly rely on positional encoding to capture the structural information Vaswani et al.
+(2017); Dosovitskiy et al.
+(2021); Su et al.
+(2024).
 Designing graph positional encoding is challenging compared to the counterpart in Euclidean spaces, due to the irregular structure and the symmetry to permutation.
-Widely used graph positional encoding includes absolute PE: LapPE Dwivedi and Bresson, (2021); Huang et al., (2024); and relative PE: shortest-path distance Ying et al., (2021), resistance-distance Zhang et al., 2023b , and RRWP Ma et al., (2023).
-Recent works Black et al., (2024); Zhang et al., 2024b study the connections between absolute PE and relative PE.
-Besides the aforementioned PE, there exist several structural encoding (SE) approaches that aim to enhance MPNNs, e.g., RWSE Dwivedi et al., 2022b , substructure counting Bouritsas et al., (2022), and homomorphism counting Jin et al., (2024).
+Widely used graph positional encoding includes absolute PE: LapPE Dwivedi and Bresson (2021); Huang et al.
+(2024); and relative PE: shortest-path distance Ying et al.
+(2021), resistance-distance Zhang et al.
+(2023b), and RRWP Ma et al.
+(2023).
+Recent works Black et al.
+(2024); Zhang et al.
+(2024b) study the connections between absolute PE and relative PE.
+Besides the aforementioned PE, there exist several structural encoding (SE) approaches that aim to enhance MPNNs, e.g., RWSE Dwivedi et al.
+(2022b), substructure counting Bouritsas et al.
+(2022), and homomorphism counting Jin et al.
+(2024).
 These SEs can effectively improve the empirical performance and/or theoretical expressivity of MPNNs.
@@
 D.4 Higher-order GNNs
-Besides graph Transformers, inspired by the $K$-WL Weisfeiler and Leman, (1968) and $K$-Folklore WL Cai et al., (1992) frameworks, $K$-GNNs Morris et al., (2019); Zhou et al., (2023); Maron et al., (2019) uplift GNNs to higher-order, treating a tuple of $K$ nodes as a token and adapting the color refinement algorithms accordingly.
-Some other higher-order GNNs Bodnar et al., (2022, 2021) are less closely related to $K$-WL but still perform well.
+Besides graph Transformers, inspired by the $K$-WL Weisfeiler and Leman (1968) and $K$-Folklore WL Cai et al.
+(1992) frameworks, $K$-GNNs Morris et al.
+(2019); Zhou et al.
+(2023); Maron et al.
+(2019) uplift GNNs to higher-order, treating a tuple of $K$ nodes as a token and adapting the color refinement algorithms accordingly.
+Some other higher-order GNNs Bodnar et al.
+(2022; 2021) are less closely related to $K$-WL but still perform well.
 $K$-GNNs can reach theoretical expressivity bounded by $K$-WL, but are typically computationally costly, with $O(N^{K})$ computational complexity.
 D.5 Subgraph GNNs
-Subgraph GNNs Bevilacqua et al., (2022); Frasca et al., (2022); Zhang et al., 2024a ; Zhang and Li, (2021); Huang et al., (2023) focus on improving the expressivity of GNNs beyond 1-WL.
+Subgraph GNNs Bevilacqua et al.
+(2022); Frasca et al.
+(2022); Zhang et al.
+(2024a); Zhang and Li (2021); Huang et al.
+(2023) focus on improving the expressivity of GNNs beyond 1-WL.
 Unlike graph Transformers and higher-order GNNs, subgraph GNNs typically do not change the model architecture.
 Instead, MPNNs are trained using a novel learning pipeline: a graph is split into multiple subgraphs, subgraph representations are learned independently, and then the subgraph representations are merged to obtain the graph representation.
-Node-level subgraph GNNs are typically bounded by $3$-WL Frasca et al., (2022), while the expressivity of edge-level subgraph GNNs is still under-explored.
-Huang et al., (2023) reveal that the expressivity of edge-level GNNs might partially go beyond $3$-WL.
+Node-level subgraph GNNs are typically bounded by $3$-WL Frasca et al.
+(2022), while the expressivity of edge-level subgraph GNNs is still under-explored.
+Huang et al.
+(2023) reveals that the expressivity of edge-level GNNs might partially go beyond $3$-WL.
 Note that subgraph GNNs usually have high memory requirements due to the need to save multiple copies for each input graph.
@@
 4 and Appx.
-E.3 of Zhang et al., 2023b , as well as in Sec.
-5 of Zhang et al., 2024b .
+E.3 of Zhang et al.
+(2023b), as well as in Sec.
+5 of Zhang et al.
+(2024b).
 The discussion on RRWP can be found in Sec.
-3.1.1 of Ma et al., (2023).
+3.1.1 of Ma et al.
+(2023).
 The empirical study of realized expressivity (in Sec.
 5.1) also matches the theoretical expressivity of PPGT.
-Based on the GD-WL analysis framework Zhang et al., 2023b , the proof is straightforward. However, we still provide a simple proof for the completeness of the conclusion.
+Based on the GD-WL analysis framework Zhang et al. (2023b), the proof is straightforward. However, we still provide a simple proof for the completeness of the conclusion.
 Proposition E.1.
@@
 Lemma E.2.
-(Lemma 5 of  Xu et al., (2019)) For any countable set $\mathcal{X}$, there exists a function $f:\mathcal{X}\to\mathbb{R}^{n}$ such that $h(\hat{\mathcal{X}}):=\sum_{x\in\hat{\mathcal{X}}}f(x)$ is unique for each multiset $\hat{\mathcal{X}}\in\mathcal{X}$ of bounded size.
+(Lemma 5 of  Xu et al.
+(2019)) For any countable set $\mathcal{X}$, there exists a function $f:\mathcal{X}\to\mathbb{R}^{n}$ such that $h(\hat{\mathcal{X}}):=\sum_{x\in\hat{\mathcal{X}}}f(x)$ is unique for each multiset $\hat{\mathcal{X}}\in\mathcal{X}$ of bounded size.
 Moreover, for some function $\phi$, any multiset function $g$ can be decomposed as $g(\hat{\mathcal{X}})=\phi(\sum_{x\in\hat{\mathcal{X}}}f(x))$.
 Proof of Proposition E.1.
-In this proof, we consider shortest-path distance (SPD) as an example of generalized distance (GD), denoted as $\text{d}_{G}^{\text{SPD}}$, which can be directly extended to other GDs such as the resistance distance (RD) Zhang et al., 2023b and RRWP Ma et al., (2023).
+In this proof, we consider shortest-path distance (SPD) as an example of generalized distance (GD), denoted as $\text{d}_{G}^{\text{SPD}}$, which can be directly extended to other GDs such as the resistance distance (RD) Zhang et al.
+(2023b) and RRWP Ma et al.
+(2023).
 Note that the choice of GD determines the practical expressiveness of GD-WL.
@@
 We note that since $\mathcal{D}_{n}$ is a finite set, its elements can be listed as $\mathcal{D}_{n}=\{d_{G,1},\cdots,d_{G,|\mathcal{D}_{n}|}\}$.
-Then the GD-WL aggregation at the $\ell$-th iteration in Eq. (13) can be equivalently rewritten as (See Theorem E.3 in Zhang et al., 2023b ):
+Then the GD-WL aggregation at the $\ell$-th iteration in Eq. (13) can be equivalently rewritten as (See Theorem E.3 in Zhang et al. (2023b)):
 $\chi^{\ell}_{G}(v):=\text{hash}\Big(\chi_{G}^{\ell,1}(v),\chi_{G}^{\ell,2}(v),\cdots,\chi_{G}^{\ell,|\mathcal{D}_{n}|}(v)\Big)\,$,
@@
 We construct the $k$-th multiset by injectively aggregating the colors of all nodes $u\in\mathcal{V}$ at a specific distance $d_{G,k}$ from node $v$.
-Assuming the color of each node $\chi^{l}_{G}(v)$ is represented as a vector $\mathbf{x}_{v}^{(l)}\in\mathbb{R}^{C}$, by setting the query and key projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K}$ as zero matrices and $\theta$ as a zero function following Zhang et al., 2023b , the attention layer of PPGT with URPE of the $h$-th head (Eq.
-5) can be written as
+Assuming the color of each node $\chi^{l}_{G}(v)$ is represented as a vector $\mathbf{x}_{v}^{(l)}\in\mathbb{R}^{C}$, by setting the query and key projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K}$ as zero matrices and $\theta$ as a zero function following Zhang et al.
+(2023b), the attention layer of PPGT with URPE of the $h$-th head (Eq.
+(5)) can be written as
 $\hat{\mathbf{x}}^{(l),h}_{v}:=\frac{1}{|\mathcal{V}|}\sum_{u\in\mathcal{V}}(\mathbf{W}_{O}^{h}\mathbf{W}_{V}^{h}\mathbf{x}^{(l)}_{u})\cdot\phi^{h}\big(\text{d}_{G}(u,v)\big)\,.$ (16)
@@
 Therefore, with a sufficient number of attention heads and a sufficient number of layers, PPGT is as powerful as GD-WL in distinguishing non-isomorphic graphs, which concludes the proof. ∎
-It is worth mentioning that the expressivity upper bound of GD-WL (i.e., 2-FWL/3-WL), given by Theorem 4.5 in Zhang et al., 2023b , is based on the usage of generalized distance (e.g., resistance distance) as graph PE.
+It is worth mentioning that the expressivity upper bound of GD-WL (i.e., 2-FWL/3-WL), given by Theorem 4.5 in Zhang et al.
+(2023b), is based on the usage of generalized distance (e.g., resistance distance) as graph PE.
 If we add PE that provides finer structural information, PPGT might surpass the 3-WL expressivity upper bound.
@@
 From Eq. (1), we immediately have:
-$\text{RMSN}(c\cdot\mathbf{x})$ $:=\frac{c\cdot\mathbf{x}}{\frac{1}{\sqrt{D}}\|c\cdot\mathbf{x}\|}\cdot\boldsymbol{\gamma}\,,$
+$\text{RMSN}(c\cdot\mathbf{x})$ $:=\frac{c\cdot\mathbf{x}}{\frac{1}{\sqrt{D}}\|c\cdot\mathbf{x}\|}\cdot\bm{\gamma}\,$,
 […]
@@
 Appendix F Limitations and Discussion
-As a plain Transformers, PPGTs still require $O(N^{2})$ computational complexity like other plain Transformers.
+As plain Transformers, PPGTs still require $O(N^{2})$ computational complexity like other plain Transformers.
 This limits the ability of PPGTs to handle very large-size inputs, which are known as large-scale graphs in the context of graph learning.
@@
 This type of dataset is the usual scenario for training large foundation models.
-  • Large-scale-graph dataset (many nodes in a graph): This corresponds to the challenges in long-context learning in language tasks and high-resolution image processing in vision tasks.
+  • Large-scale-graph dataset (many nodes in a graph): This corresponds to the challenges in long-context learning in language tasks and gigapixel image processing in vision tasks.
 The challenges mainly lie in the computational efficiency and memory consumption of the models.
@@
 As plain Transformers, PPGTs can potentially adopt several efficiency techniques developed for plain Transformers in graph-related and other domains, directly or with modifications.
-BigBird Zaheer et al., (2020) and Exphormer Shirzad et al., (2023) propose building sparse attention mechanisms through sparse computational graphs, using random graphs and expander graphs for language tasks and graph learning tasks, respectively.
+BigBird Zaheer et al.
+(2020) and Exphormer Shirzad et al.
+(2023) propose building sparse attention mechanisms through sparse computational graphs, using random graphs and expander graphs for language tasks and graph learning tasks, respectively.
 They can be directly adopted by PPGTs.
 […]
-Ainslie et al., (2023) propose grouped-query attention (GQA) to save the memory bandwidth.
-It is a more specific design for decoder-only Transformers (for KV-cache), and might not introduce remarkable benefits to encoder-based Transformers, such as ViTs Dosovitskiy et al., (2021); Touvron et al., 2021a ; Liu et al., (2021), Time-series Transformers Nie et al., (2023); Zhang et al., 2024c and most GTs Kreuzer et al., (2021); Ying et al., (2021); Zhang et al., 2023b ; Ma et al., (2023) including PPGTs.
+Ainslie et al.
+(2023) proposes grouped-query attention (GQA) to save the memory bandwidth.
+It is a more specific design for decoder-only Transformers (for KV-cache), and might not introduce remarkable benefits to encoder-based Transformers, such as ViTs Dosovitskiy et al.
+(2021); Touvron et al.
+(2021a); Liu et al.
+(2021), Time-series Transformers Nie et al.
+(2023); Zhang et al.
+(2024c) and most GTs Kreuzer et al.
+(2021); Ying et al.
+(2021); Zhang et al.
+(2023b); Ma et al.
+(2023) including PPGTs.
 We mention only a few efficiency techniques for plain Transformers here.
@@
 This paper presents work whose goal is to advance the field of Geometric/Graph Deep Learning.
-There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.